### v2

# DiVerify: Hardening Identity-Based Software Signing with Programmable Diverse-Context Scopes

Abstract
Code signing enables software developers to digitally sign their code using cryptographic keys, thereby associating the code with a specific key. This key is then linked to an identity (e.g., through an identity provider), allowing users to establish trust in the origin of the signature and verify both the code’s origin and integrity. However, this code–identity binding is only as trustworthy as the mechanisms enforcing it. State-of-the-art identity-based code signing schemes have a major shortcoming: they fail to provide verifiable information about the context in which a signature is generated. If an identity verification server is compromised or the signing client behaves maliciously, the resulting signature may falsely suggest a trustworthy origin, despite the absence of actual developer intent.
To address these issues, we propose a diverse identity verification approach that reduces reliance on a single source of verification and enforces stronger guarantees around the signing process itself. By combining multiple identity signals with verifiable execution environments, our system improves confidence that signatures reflect the intent of a legitimate user, produced under expected conditions. Signing in our DiVerify prototype incurs only a few kilobytes of additional storage (less than 0.4% of the average package size in widely used ecosystems like PyPI) In DiVerify, signing completes in under 100ms on a typical deployment.
“Yo soy yo y mi circunstancia; y si no la salvo a ella no me salvo yo.”111“I am I and my circumstance; and, if I do not save it, I do not save myself.” — José Ortega y Gasset
1 Introduction
Software supply chain attacks have become a concern due to their large-scale impact on software users’ safety [73, 71, 9] These attacks involve the injection of malicious code into a software artifact and subsequent exploitation in downstream systems [64]. In response, government and industry groups have begun to advocate for improved provenance along the software supply chain [13, 26, 20].
The most common provenance method for software supply chains is software signing. Software providers attach a digital signature to an artifact using public key cryptography, and software consumers can then test this signature to detect tampering [14]. Despite decades of use [74, 19], adoption has been limited by usability [35] and key management challenges [79, 36]. To address this, identity-based signing platforms emerged [78], still grounded in public key cryptography [62, 49], but linking signatures to verified identities via identity providers (IdPs) and short-lived ephemeral keys to allow a single signing event.
However, the state-of-the-art identity-based signing solutions are not resilient. If any party involved in the signing operation is compromised, then the provenance guarantees are lost. This limitation manifests in two ways. First, identity-based signing systems rely on a single IdP to identify a signer and issue authentication tokens used for signing. If this identity provider is compromised, it renders all associated signing events untrustworthy. Many recent attacks have targeted identity providers, including the Midnight Blizzard intrusion [55], Apple’s SSO [6], and attacks on Facebook [69, 76].
The second limitation is the signing client, which brokers interactions between the software author, the identity provider, and the repository. These clients are entrusted with signing capabilities for arbitrary code during the validity of their ephemeral key. In this case, the security of code signing on the client side relies on the effectiveness of client-side security measures. If compromised, they can sign malicious software that appears fully legitimate. An example is the attack by the North Korea-based threat actor Diamond Sleet (ZINC) [54], where attackers exploited a code-signing system’s privileges to sign malicious variants of an application developed by CyberLink, making it appear legitimate. Current solutions rely on retrospective monitoring of signature logs to detect unauthorized signing [62], which only occurs once malicious software has been distributed and utilized.
These two problems share a common cause: inadequate transparency and verifiability of the conditions under which a signature was produced. Current systems provide no mechanisms for verifiers to assess the signer’s security posture or signing context, creating a blind spot that attackers can exploit using valid credentials under malicious conditions. To address this, software signing systems must evolve beyond verifying who signed a payload, and must also verifiably communicate how a signature was produced – under what conditions, on what systems, and with what constraints. In other words, we need mechanisms to make the signing context itself verifiable, so that trust is no longer concentrated in a static identity or a privileged client but distributed across a transparent, auditable chain of evidence.
Towards this goal, in this paper we present DiVerify, a system that enhances identity-based signing by enforcing a verifiable, contextual security posture at signing time. DiVerify introduces a programmable scope verification framework that expands the notion of signer identity to include a flexible set of verifiable attributes – called DiVerify scopes. DiVerify operates under a zero-trust assumption toward signing clients. Rather than assuming the integrity of the client environment, it requires that all sensitive signing operations occur within a hardware-protected Trusted Execution Environment (TEE). Each signing operation produces a DiVerify proof: a cryptographically verifiable attestation of both identity attributes and the runtime integrity of the signing environment. This enables the enforcement of fine-grained, context-aware policies for signature verification, reducing implicit trust and raising the bar for a successful client compromise.
The core novelty of DiVerify is that it extends identity-based code signing to enforce how software is signed, not just by whom. DiVerify enables the enforcement of multi-independent identity and client attestations at signing time, ensuring verifiable secure context for software signatures.
We evaluate DiVerify in terms of its security guarantees and operational overhead. To assess security, we show how different DiVerify instantiations can balance security and performance across different variants to ease integration into existing identity-based signing systems, and highlight how they mitigate the adversaries described in our threat model. To assess overhead, we quantify the performance and storage costs of signing and verification operations across the different variants. To demonstrate its practical relevance, we analyze several historical software supply chain attacks and illustrate how DiVerify’s model could have helped detect or prevent them. Our evaluation shows that DiVerify introduces minimal overhead compared to current identity-based software signing solutions with signing operations complete in under 100ms for regular deployment, and verification adds only  11% latency compared to legacy-compatible approaches, highlighting its feasibility for real-world adoption.
In summary, we make the following contributions:
  1. 1.
    A novel framework for context-aware identity verification in code signing, which, for the first time, combines multiple identity proofs with hardware attestation.
  2. 2.
    Design and implementation of DiVerify, a prototype that enforces context-bound signing policies using TEE-based attestations, an approach not present in prior signing tools.
  3. 3.
    Evaluation of security and performance under different settings, demonstrating that enforcing multi-identity verification with TEE-backed signing is practical, which was previously unexamined.
  4. 4.
    Security analysis showing that DiVerify would have prevented several real-world supply-chain attacks that prevailing solutions could not have.
2 Background
To describe DiVerify, we first discuss its context (i.e., the software supply chain) and its role (i.e., code signing). The DiVerify protocol also requires understanding authentication mechanisms used to establish identity and context, and remote attestation, which we leverage in our implementation.
2.1 Software Supply Chain and Code Signing
The software supply chain involves all the activities carried out in the task of creating and distributing software, such as coding, compiling, testing, packaging and deployment [79, 64]. These activities rely on a wide range of tools, libraries, frameworks, and services creating a complex web of dependencies [42, 83][42, 83]. This complexity makes it a prime target for attackers, as seen in incidents like SolarWinds [24] and Log4j [60]. As a result, organizations are increasingly adopting practices like code signing, improved dependency management and vetting [68, 18], continuous monitoring [84], zero-trust architectures [46, 86, 1], and Software Bills of Materials (SBOM) [61] to ensure transparency, integrity, and compartmentalization throughout the supply chain.
Code signing allows a verifier to establish trust with a software producer. The process uses a keypair $(k_{p},k_{s})$ and an artifact $(SW_{a})$ to generate a signature $(S_{sw})$. A verifier will then use $(K_{p},SW_{a},S_{sw})$ to verify whether $(SW_{a})$ is, first, integral (i.e., hasn’t been tampered with), and second, belongs to a trusted party. Traditional tools like OpenPGP [67] use manual key management and user-controlled trust models [8, 87], requiring developers to generate, safeguard, and distribute long-lived keys, introducing usability challenges and risks of compromise. Identity-based solutions such as Sigstore [62], OpenPubKey [25], and Keyfactor [38] simplify this by using ephemeral keys and identity federation protocols (e.g., OIDC, OAuth, SAML). These allow developers to prove identity ownership via standard authentication flows, then linking identities to signing keys through certificates or embedded tokens.
2.2 Authentication Mechanisms
Identity-based signing tool uses authentication to verify the identity of a signer. Authentication mechanisms vary, but generally fall into categories based on the type of claim they make [27, 51]: identity, possession, or context.
Identity claims verify who a signer is, using protocols like OAuth 2.0 [63] and OpenID Connect [65]. These protocols issue signed tokens (e.g., JWTs [4]) from trusted providers like Google. Distributed Identities (DiD) [89] and CryptID [28] provide mechanisms similar to OIDC but act as a meta-protocol enabling identity-specific claims.
Possession claims prove control over physical elements such as a hardware token, smartcard, or cryptographic key. The authenticated party demonstrates possession of a secret (e.g., a private key) through a cryptographic challenge-response protocol. Protocols like FIDO2/WebAuthn, use devices (e.g., YubiKeys [91]) to protect the private key and perform challenge responses locally ensuring that only someone with the device can authenticate.
Contextual claims assert properties of the environment in which events occur, focusing on system state rather than identity or key possession. They support trust decisions based on runtime conditions like device or software integrity. Techniques include device fingerprinting [75], TPM-based measurements [53, 32], and kernel-level attestation via Linux’s Integrity Measurement Architecture (IMA) [48]. In DiVerify, such claims help bind events to execution contexts.
2.3 Trusted Execution Environments
Trusted Execution Environments (TEEs) isolate processes using hardware mechanisms like secure memory and privileged execution contexts [3]. Remote Attestation enables verifiers (e.g., a remote user or system) to confirm software execution within a secure and trusted environment on a remote machine. This is achieved through a cryptographic protocol in which the remote system generates a signed statement known as an attestation report or quote that includes a measurement (e.g., hash) of the software and configuration loaded into a secure execution environment. The verifier can compare this measurement against a known trusted value to determine whether the software executing is indeed the expected one.
Trust in remote attestation relies on hardware-rooted chains of trust, such as those provided by Intel Software Guard Extensions (SGX) [17], Trusted Domain Extensions (TDX) [31] which include embedded attestation keys and associated endorsement certificates issued by the hardware manufacturer. These cryptographic materials ensure that only legitimate, manufacturer-approved hardware can produce valid attestation signatures.
3 Parties, Roles, and Threat Model
Next we formalize identity-based software signing systems, the context in which they operate, and the security goals of both identity-based software signing in general, and DiVerify in particular.
3.1 Parties and Roles
We consider a standard identity-based software signing setting with the following parties interacting with each other:
Signer: an entity (i.e., a developer or an automated system) who authored the software artifact and intends to generate a signature $(S_{s}w)$ for it.
Client: a tool or service with which the Signer interacts to produce such signatures. To do so, a Client needs to know the Signer’s identity to complete a signing request. This request involves generating key material and interacting with an Identity Provider to establish trust in such material.
Identity Provider (IdP): an entity that authenticates the Signer through the Client and provides qualified attributes about the User. In DiVerify we rely on a general-purpose instantiation of an IdP we call scope providers as described in section˜5.3.
Package repository: a storage location where software packages and their signatures are published, shared, and retrieved for distribution and verification, e.g., NPM or Hugging Face.
Verifier: The counterpart of the Signer. It verifies the signed software to establish trust by validating a signature against a software artifact using a certificate or public key.
Figure 1: State-of-art identity-based code signing system model, highlighting the types of attacker capabilities that such systems are not resilient against. These attack types are discussed in section˜3.2.
A typical high-level interaction flow is illustrated in Figure˜1. There are five steps. 1. The signer initiates a signing request by interacting with a client. 2. The client redirects her to an IdP to authenticate. 3. The signer authenticates, an identity token is obtained, and sent to the client. 4. Optionally, the client requests a signing certificate (which binds the identity to a signing key) from a certificate authority (or CA). 5. Finally, the client signs the software and uploads the signature to a repository.
A successful flow allows a verifier to obtain a $(S_{sw},K_{p},SW_{a})$ tuple from a source (e.g., a repository) and ensure the artifact is not tampered with.
3.2 Threat Model
Attackers intend to subvert the flow described above to deliver tampered software to verifiers (i.e., software users). In order to do so, they will impersonate, tamper with, and/or attack all the roles described above. A successful attack is one that causes a verifier to accept a malicious artifact $(SW_{a})$.
At a fundamental level, we want to ensure code signatures accurately link to the legitimate Signer’s identity and reflect their intent. Hence, we consider threats where adversaries seek to produce valid signatures without the consent of the legitimate identity owner.
Attacker Capabilities
We describe three main types of adversaries with the following capabilities:
  1. 1.
    Type I: Credential Compromise. The attacker compromises a signer’s account, typically through credential theft (e.g., phishing, password leaks, session hijack). Once the attacker has access to the account, they can use it to authenticate and sign software on behalf of the signer without their consent. Type I attackers are successful if they are able to identify themselves on behalf of the Signer against an IdP.
  2. 2.
    Type II: Compromised or Rogue Identity Provider. The IdP is compromised or malicious, and can issue identity tokens for a signer who did not authenticate. An attacker can then use the fraudulently issued token to impersonate the signer gaining signing access. Type II attackers are able to impersonate an IdP and thus a Signer without interaction with either.
  3. 3.
    Type III: Compromised Signing Tool. The attacker can tamper with the client software to sign software beyond what the user originally intended, allowing unauthorized code to be signed.
Examples of each type
Recent events illustrate:
  1. 1.
    ESLint (Type I: Credential Compromise): ESLint, a widely used JavaScript linter, was compromised in 2018 when attackers used stolen npm credentials to upload malicious package [92]. Because the system relied only on token-based authentication, it could not tell legitimate actions from malicious ones.
  2. 2.
    Midnight Blizzard (Type II: IdP Compromise): In 2023, the nation-state actor Midnight Blizzard compromised a legacy Microsoft test account to create malicious OAuth apps with elevated privileges [55]. These apps granted persistent access to Exchange Online, bypassing identity checks. Since the tokens were issued by a trusted provider and used valid credentials, the system lacked the context to detect the compromise.
  3. 3.
    Codecov (Type III: Tool Compromise): A widely used code coverage tool integrated into CI was compromised in 2021, when attackers added code to exfiltrate sensitive information [10]. The attack persisted undetected as users implicitly trusted the tool, with no way to verify its runtime state. Other previous attacks have also taken this form of abuse [85, 40, 44].
Assumptions
We assume legitimate signers do not sign untrusted code, verifiers properly check all signatures, the TEE setup remains secure, and not all identity providers can be compromised at the same time. Rationales and additional assumptions are discussed in Appendix˜A.
4 System Goals and Overview
This section shows our security and deployment goals and gives an overview of DiVerify’s approach.
4.1 System Goals
In order to address the attackers described in our Threat Model, DiVerify aims to extend authentication beyond traditional credentials; harden the security posture of signing clients; and limit reliance on any single identity provider by requiring multiple, independent sources of verification. We formalize these notions into goals:
Security Goals:
  1. S1
    Resilience Against Identity Theft: Signing events are allowed only after verification from multiple independent sources, which limits trust assumptions on identity providers. This way, if an attacker compromises a Signer’s credentials, an attacker is unable to complete a signer request without a collection of these.
  2. S2
    Resilience Against IdP compromise: In addition to diversity of credentials, a Signer can request multiple identity providers to prove a diverse identity claim. This ensures that a compromised provider cannot impersonate its user unless it has also compromised a threshold number of other independent providers.
  3. S3
    Resilience Against Compromised Clients: It should be possible to cryptographically verify that a signature was generated by a trusted tool and not a compromised version. This is possible by having signing information include information about the signing context as well as identity claims.
Deployment Goals:
In addition to the security goals, DiVerify is also designed to facilitate adoption within existing software supply chain tools:
  1. D1
    Backwards compatibility in a mixed model within the code signing ecosystems
  2. D2
    Little-to-no changes to existing tooling
4.2 System Overview
DiVerify follows a disaggregated trust model where trust in a signature is subject to a combination of independently verifiable claims about identity, context, and delegated trust. In practice, this manifests in a similar way as the identity-based software signing flow described above. However, in contrast, we rely on a generalized party, a scope provider, which replaces the role of an IdP. In order to dis-aggregate trust, a set of Scope Providers will provide individual proofs regarding the Signer’s properties.
Figure 2 presents a high-level view of the system. In the same way as a regular identity-based software signing setting, the flow initiates when a Signer requests that a client performs the signing process. Instead of requesting an identity token from an IdP, it will query a set of scope providers to obtain multiple scopes. These scopes can include, e.g., OIDC identity token as proof of identity, a device fingerprint or OS attestation as proof of environment, and a hardware security token proof as a second factor. In parallel, the client also collects integrity evidence about the signing environment itself.
Having collected these scopes, the client must aggregate them into a DiVerify proof which shows how the signer was authenticated, and also embed a proof of its own integrity. This can be carried out in various ways such as using TPM-ME and a seal/unseal system [30], or by moving a portion of the client code into a separate party inside a TEE. The resulting DiVerify proof serves a role similar to an identity-based signing certificate, but with richer information that a verifier can use to assess whether they trust a particular signature over a piece of software. Lastly, the verifier is responsible for validating these proofs and verifying the resulting signatures.
Figure 2: DiVerify Overview. DiVerify extends traditional identity providers into scope providers, which assert claims about the signer. The DiVerify daemon, running in a secure enclave, coordinates and proves its own integrity to a verifier.
For example, consider a flow where a signer Alice wants to sign an artifact in a DiVerify setting. They want verifiers to know that not only are they alice@software.sh, but also that they signed their software inside her MacBook Pro (identified with the fingerprint 0xCAFEC0FFEE), and using a second factor from her Yubikey. Users of Alice’s software are aware of this, and will use a DiVerify policy to agree that valid software from Alice needs to contain proofs of the above.
Alice will then follow this flow to generate the signature, which Bob can then validate:
  1. 1.
    Alice will issue a signing request through her client tool, specifying a trust level (i.e., the set of properties described above).
  2. 2.
    The client acting on behalf of the Alice prepares a signing request object, determines which scopes meet the trust level and prompts the her to authenticate with the corresponding scope providers.
  3. 3.
    For each scope requested, each provider will issue signed scope to the client if Alice is successful. In this particular case, Alice authenticates against an identity provider to prove she is indeed alice@software.sh, she obtains a fingerprint from her operating system non-interactively, and touches her Yubikey to generate a second factor scope signature.
  4. 4.
    The client verifies the returned scopes, signs the software, and appends a last integrity check to prove its own integrity (e.g., by creating a TEE remote-attestation object). Then it aggregates these scopes and the attestation into a DiVerify proof.
  5. 5.
    The signed artifact, along with its DiVerify proof, is submitted to a package repository for storage and distribution.
  6. 6.
    Bob retrieves the artifact, extracts its associated policy, and checks that: (a) the signature is valid and it applies to the obtained artifact. (b) the provided proof satisfies the agreed upon policy (that describes which scopes are needed and their expected values). (c) the proof was generated by a legitimate and trustworthy client software.
[…]
Of note, some of the techniques of DiVerify overlap. For example, the use of contextual scopes will also reduce the likelihood that an attacker is able to sign on behalf of the client using stolen credentials. This is because an attacker must also be able to initiate a signing request in the same context (e.g., hardware) as the Signer.
5 DiVerify Architecture
Having provided an overview of DiVerify , we move on to describe the components of DiVerify , followed by the protocol of how they interact to support secure and verifiable signing.
5.1 Components
Scope Providers
manage and issue claims about a signer’s identity, attributes, or environment context. Each operates within a distinct trust domain, that is, an independently managed authority responsible for issuing and validating its own set of claims about signers. They follow a specific protocol to verify properties such as control of an email address, possession of a cryptographic key, or ownership of a device and then issues a signed statement attesting to that property. DiVerify does not propose new protocols for scope providers, but builds on existing authentication standards whenever possible. Scope providers must be able to authenticate themselves to the tool and prevent scope replay to ensure that claims are both trustworthy and fresh. Since not all providers support both of these properties natively, DiVerify introduces strategies to “wrap” them and implement arbitrary scope providers at a meta-protocol level. We discuss these strategies in section˜5.3.
Client
is a signer-facing component responsible for initiating signing requests. This can be instantiated as a command-line tool or library integrated into developer environments. The client operates within the signer’s local context and has access to the artifact to sign. It serves as the point of integration between the signer and the rest of the DiVerify architecture. Because the client runs in a potentially untrusted setting, we split the functionality of a traditional signing client into two parts: a lightweight front-end Client and a privileged Daemon.
DiVerify Daemon
is a long-running process responsible for coordinating with scope providers and performing signing operations. Importantly, it runs inside a TEE on the user’s machine. The daemon thus serves as a secure execution environment that can verify scopes and construct proofs with a higher level of assurance, being isolated from the rest of the system. It can also generate a remote attestation to prove its own integrity to external verifiers.
Verifier
is an end-user side component responsible for validating signed artifacts and assessing their trustworthiness. It operates independently of the signing process and has no privileged relationship with the signer. A DiVerify verifier will receive a triple $(K_{p},SW_{a},S_{sw})$ as well as a policy $P_{sw}$ and return either $OK$ if the software is trusted or $ERROR$ if any of the validation steps failed.
5.2 DiVerify Protocol
5.2.1 Signing Request Initialization
DiVerify-enabled signing tools can advertise which scope providers they support, allowing signers to know what factors can be used. We define a trust levels configuration that groups scopes into levels based on the types of claims. Higher trust levels correspond to requiring more diverse claims (an example is shown in Listing LABEL:lst:diverify-config-structure in the Appendix), increasing overall confidence in the signer’s legitimacy. When the DiVerify daemon starts, it establishes channels to interact with the client and with scope providers.
To initialize, a Signer interacts with a Client to initialize a signing request. This can be done by issuing a command through the DiVerify client. The parameters to initialize a signing request include trust level, and an identifier (e.g., a filename, or docker image tag) of the software artifact to sign.
5.2.2 Scope Retrieval and Signing
With a signing request initialized, the client submits a request to the daemon. Following Algorithm 1, the daemon first determines which scopes are required for the requested trust level (line 2). These required scopes are independent, and can be obtained in parallel from their respective providers. The daemon generates a fresh nonce to use in all scope requests to bind them to this signing session. It then iterates over each target scope provider and initiates an authentication protocol to obtain the scope (lines 4–7).
Algorithm 1 DiVerify Daemon Flow to Sign Payload
1:In: Artifact AA, TrustLevel $L$
2:Requires: Scope Provider $P$, TrustConfig
3:Out: ArtifactSignature $S$, DiVerify Proof dvpdvp
4:procedure SignRequest(AA, $L$)
5:  $req\_scopes\leftarrow\text{TrustConfig}(L)$
6:  $scopes\leftarrow\{\}$
7:  for $m\in req\_scopes$ do
8:   $r\leftarrow\text{AuthenticateSigner}({P}_{m})$
9:   $scopes[m]\leftarrow r$
10:  end for
11:  if not ValidScope(scopesscopesthen
12:   return Err
13:  end if
14:  $sk,pk=\text{GenerateKeyPair}()$
15:  $S\leftarrow\text{Sign}_{sk}(A)$
16:  $custom\_data\leftarrow\text{Sign}_{sk}(hash(scopes))$
17:  $quote\leftarrow\text{GetQuote}(custom\_data)$
18:  $dvp\leftarrow\text{MakeDiVerifyProof}(scopes,quote,pk)$
19:  return $S$, dvpdvp
20:end procedure
Each request follow a DiVerify specific meta-protocol. The protocol proceeds as follows: first, the daemon submits a request such as $request(n,claim)$. Each provider responds with a tuple $(\mathit{sign(proof)},\ \mathit{proof})$, where $\mathit{proof}$ is a protocol-specific claim (e.g., an oidc token), and the accompanying signature binds it to the daemon’s nonce. Of note, the nonce itself is embedded inside of the proofproof. This could be trivially surfaced as part of the top-level response (and turn the tuple above into a triple), but this would not retain backward compatibility to other natively supported protocols such as OIDC.
On the scope provider side, scopes may request the signer to prove the claimclaim. For example, if a Signer requests an OIDC provider to sign a scope for alice@software.sh, the provider may request a $username,password$ tuple to be input through the client or a browser. Other scope providers may not require user interaction, and will reply with the proof directly.
Once the daemon receives scope responses from providers, it verifies each tuple $(\mathit{sign(proof)},\ \mathit{proof})$ by checking that the signature is valid and that the proof includes the correct daemon-generated nonce (line 8). Since the daemon is trusted, the Signature from the scope provider does not need to be forwarded to the verifier. After successful verification, the daemon proceeds to generate ephemeral signing keys and perform the signing operation.
5.2.3 Attestation Binding
The daemon must prove to a verifier that it is running untampered inside an enclave via remote attestation. To prevent a client from replaying attestations using scopes or signing key it controls, the daemon must cryptographically bind the scope and signing key to the attestation.
To achieve this, the daemon first signs a hash of the validated scopes using the signing key (line 13). It then embeds the resulting signature into the attestation quote (line 14).
Because this field is included in the signed remote attestation quote, it cryptographically binds the signing key to the TEE identity and state.
5.2.4 DiVerify Proof Construction
Since a verifier needs to have a direct view on the scope claims, and key used for signing to allow it to verify a signed artifact, DiVerify needs to be able to provide that information. This is done by providing a DiVerify proof (as shown in Listing LABEL:lst:proof-structure in the Appendix), which consists of the following: the validated scopes (i.e., the data that the daemon validated and signed), the signing public key pair, and the remote attestation. The daemon then bundles the proof and the signature, and send to the client who uploads it to the package repository.
5.2.5 Verification
To verify a signature, a verifier needs the signed artifact, the generated signature, DiVerify proof, and policy. The DiVerify policy defines expected signer claims and a logical rule of how these claims must combine to satisfy trust. We discuss its semantics below and how it is securely distributed using TUF[72] in section˜5.4.
The verification process follows the flow outlined in Algorithm 2. The procedure begins with the verifier retrieving the current verification policy from TUF. It then extracts the attestation quote, the associated custom data, the public key, and the scopes from the DiVerify proof. Next, it validates the integrity of the quote (line 7), confirms that the scope digest matches the custom data using the public key (line 10). Then it evaluates whether the scopes comply with the retrieved policy. Finally, it validates that the artifact using the public key. If any of these checks fail, the artifact is deemed untrusted and is consequently rejected.
Algorithm 2 Artifact Verification Flow
1:Input: Artifact A, ArtifactSignature $S$, DiVerify proof dvpdvp
2:Requires: RSTUF $T$
3:Output: Boolean verification result
4:procedure VerifyArtifact(A, S)
5:  policypolicy $\leftarrow$ $T.\text{GetPolicy}()$
6:  quotequote $\leftarrow$ $dvp.\text{quote}$
7:  cutom_datacutom\_data $\leftarrow$ $quote.\text{cutom\_data}$
8:  pkpk $\leftarrow$ $dvp.\text{pubkey}$
9:  scope_digestscope\_digest $\leftarrow$ $hash(dvp.scopes)$
10:  if not ValidQuote(quotequotethen
11:   return false
12:  end if
13:  if not VerifySignature(scope_digestscope\_digest,custom_datacustom\_data,pkpkthen
14:   return false
15:  end if
16:  if not EvaluatePolicy($dvp.scopes$, policypolicythen
17:   return false
18:  end if
19:  if not VerifySignature(AA,$S$,pkpkthen
20:   return false
21:  end if
22:  return true
23:end procedure
Diverify Policy Semantics
We recognize that a legitimate signer may not always be able to satisfy all scopes at every signing instance. For example, a signer using a personal computer may authenticate via an identity provider, prove possession of a hardware token, and demonstrate that signing occurred on a specific, trusted machine fingerprint. However, when operating from a remote environment, the same signer may still authenticate using an identity provider or hardware key—but may no longer be able to prove that signing occurred on the expected, fingerprinted device.
DiVerify supports this flexibility, and allows software maintainers to define trust policies that specify which scopes are required and what constraints must hold for a signature over the software to be accepted.
In the sample policy in listing LABEL:lst:diverify-scope, each field defines an expected value that must match a corresponding claim in the DiVerify proof. The rule (lines 10) field expresses a Boolean condition over these claims, defining the minimal combination required for trust. In the sample, the policy requires the signer to authenticate with the specified identity and provider, and to present either the expected device fingerprint or a hardware token. Additionally, the proof must provide a remote attestation that matches the known trusted signer measurement.
1"identity": {
2 "email": "alice@software.sh",
3 "provider": "https://accounts.google.com",
4"device_fingerprint": {
5 "expected_fpr": "0x1234abc...",
6"signer_measurement": "sha256:5678def...",
7"security_key": {
8 "slot9a_public_key": "Fkw...EwYH",
9 "slotf9_attestation_cert": "MIE..AB",}
10"rule": "(identityANDprovider)AND(device_fingerprintORhardware_key)AND(signer_measurement)"
Listing 1: Sample DiVerify Policy Structure
5.3 Integrating Scope Providers
DiVerify requires scope providers must be able to authenticate themselves to the daemon and prevent scope replay to ensure that claims are both trustworthy and fresh. however, not all providers natively support these. To accommodate this, we introduce two strategies: wrapping and TEE integration—which support various provider capabilities. We discuss different categories of these providers and how they are supported in DiVerify.
Verifiable Providers (VP)
are able to cryptographically authenticate themselves and issue scopes that are bound to a specific context, making them verifiable and resistant to replay. For example, OIDC tokens signed by a known Identity Provider such as Google. VPs are natively supported in DiVerify with minimal changes, and transferring their scopes only requires defining a syntax for policy verification.
Static Providers (SP)
SPs do not support nonce injection, but the authenticity of the provider can still be verified through cryptographic signatures or secure communication channels. SPs are often stateless protocols that provide a fixed value whenever called. For example, this is the case of hardware keys using device attestation for PIV, where the same x509 certificate is provided every time an attestation request is made. This introduces two risks, first, without state-keeping an attacker can cache scope proofs. Second, if the value is consistent, then an attacker could also re-generate that value in another context.
We devise two strategies to support SPs:
  • Zk-based SP: the scope provider is wrapped inside a process that provides a proof of the value being a correct device fingerprint. This can be transparently verified by both the verifier and the diverify daemon, and the nonce utilized can match the nonces provided to other scope providers.
  • nonce-hash SP: in some cases, authentication is enough and its only necessary to prove the scope requested was part of a signing event. In this case the reply is hashed along an input nonce and a triple (noncenonce, measurementmeasurement, sigsig(hashhash(noncenonce|measurementmeasurement)) is returned.
Opaque Providers (OP)
OPs allow nonce injection to prevent replay attacks, but lack a reliable mechanism to verify the provider’s identity making them susceptible to spoofing if used in isolation. For example, this is the case of TPM ME and device fingerprinting writ large.
In order to support these, we introduce a broker/wrapper process that can be authenticated to provide opaque provider values. This is trivially done by wrapping existing scope provider software inside of a trusted execution environment.
Untrusted Providers (UT)
UTs support neither nonce injection nor provider verification, offering no protection against replay or impersonation attacks and representing the weakest security posture. We combine the strategies for OPs and SPs to support UTs.
For all the above, we introduce function prototypes to support the implementation of new scope providers. In general, for all providers other than VP, we utilize two strategies. The first is wrapped providers (i.e., wrapping the provider process within a trusted process). The second is TEE-based providers (i.e., transferring the provider process to a TEE).
5.4 Distribution of DiVerify Policy
DiVerify’s security relies on the tamperproof distribution of its policies, which serve as the root of trust for verification. To ensure freshness, authenticity, and integrity, we use The Update Framework (TUF) [43].
TUF uses separation of duties to isolate signing different pieces of metadata about DiVerify policies, as illustrated in Figure˜3. The idea is that while machines can be trusted to sign with online keys—or keys sitting on the repository—higher-level statements about whether there are new or updated DiVerify policies (i.e., the timestamp and snapshot metadata), only humans are allowed to sign with offline keys—or keys sitting off the repository—the policies themselves that confer trust on those packages in the first place (i.e., the root and targets metadata). This ensures that even if an attacker compromises the TUF repository, they cannot tamper with policies without being detected because they do not have access to offline keys, and so they cannot tamper with packages because their signatures would not match the expected policy. Note that this idea generalizes to different policies for different packages in a shared repository, although the exact details are left to future work.
Figure 3: DiVerify uses TUF[43] to distribute configuration and policy securely via a chain of roles. The Root role establishes trusted keys for other roles, the Targets role signs metadata about the policy, and the Snapshot and Timestamp roles ensure integrity and freshness of the repository state.
TUF requires developers to bootstrap the repository with an offline key ceremony, during which they define which keys or identities are allowed to sign which pieces of metadata (notably about DiVerify policies). After this one-time setup, developers need only sign and publish new TUF targets metadata about new DiVerify policies, and machines can be safely trusted to sign and publish only the “news” whether there are new policies at all.
When verifying a package signature, DiVerify first uses TUF to fetch and verify the latest policy. This includes validating the full signature chain, starting from the root metadata down to the targets metadata that points to the exact policy. Only after this verification succeeds does DiVerify apply the now-verified policy to evaluate the package’s signing proofs.
6 Adapting DiVerify to Existing Systems
(a) Legacy-Compatible Model
(b) Trusted Authentication Model
Figure 4: Two deployment modes for DiVerify. (a) Diverse verification without client integrity attestation. (b) Diverse verification with client integrity attestation. The latter provides stronger security by ensuring that verification results cannot be tampered with by a compromised client.
While DiVerify defines a core trust model for diverse identity verification, tool, and process integrity, the architecture admits multiple realizations with varying security guarantees. DiVerify can be incrementally adapted to fit into existing identity-based signing systems by adjusting trust assumptions, replacing components, or selectively omitting features like attestation or trusted hardware.
This section outlines how DiVerify can be instantiated in different environments, each offering tradeoff between security and deployability. These variants demonstrate how DiVerify’s underlying mechanisms can integrate with real-world existing signing solutions that operate under different constraints.
6.1 Legacy-Compatible
The legacy-compatible variant enables stronger security guarantees in existing signing workflows with minimal disruption, at the cost of a weaker posture against client compromise.
Current solutions typically rely on a single authentication mechanism - an OIDC identity token. In this variant as shown in Figure˜4a, DiVerify operates without a secure daemon. Instead, the client is trusted to honestly retrieve and present authentication claims, and the DiVerify proof becomes a collection of such claims, embedded directly into an existing root of trust (e.g., a signing certificate).
For example, in systems like Sigstore, which issues ephemeral certificates tied to OIDC identities and signing keys, the certificate could be extended to include a DiVerify proof that contains additional scopes. Similarly, in OpenPubkey, which binds signing keys to federated identity claims, the OIDC scope becomes one component in a more expressive interpretations of signer legitimacy.
Even in this relaxed model, software maintainers can define DiVerify policies that go beyond a single identity source, and constraint for such identities to sign.
However, this setting sacrifices client-side integrity guarantees. Because there is no trusted daemon or attestation, the system cannot validate client integrity. A compromised client could spoof or replay scopes. While a certificate authority may verify the authenticity of claims against the issuing scope providers, this only holds for secure and verifiable providers that are able to authenticate themselves.
6.2 Trusted Authentication
The Trusted Authentication variant (depicted in  Figure˜4b) is similar to the core DiVerify model but differs in when the daemon’s integrity is verified and who performs the verification.
As in the core model, the daemon retrieves scopes from scope providers. However, instead of proceeding to signing after collecting and validating scopes, the daemon defers signing until an external verifier attests to its integrity. This additional check helps prevent a compromised daemon from producing valid signatures.
A common choice for this external verifier is a Certificate Authority, which can embed the result of the daemon’s attestation into a signing certificate. This approach aligns with existing infrastructures such as those described in [62], where CAs verify the identity of a signer before issuing a certificate. However, the external verifier does not need to be limited to traditional CAs. Alternatives include federated identity systems, policy engines, or custom services that can enforce integrity checks before authorizing a signing operation.
By requiring the daemon’s integrity to be confirmed prior to signing, this variant provides a proactive safeguard against signature generation by a compromised daemon. However, this setting still relies on a trusted CA deployment, which introduces a single point of failure.
7 Security Evaluation
7.1 Security Analysis
[…]
Compromise Without With DiVerify
Type Diverify
Legacy Trusted Core
Compat. Auth.
[…]
Client $Circle \CIRCLE \CIRCLE$
[…]
Certificate Authority $Circle \CIRCLE -$
We consider three attack scenarios in identity-based signing, derived from our threat model (section˜3.2) and show how our core, legacy-compatible section˜6.1 and trusted authentication section˜6.2 DiVerify variant protects against them.
First, the compromise of the scope provider to issue a fraudulent signer scope; second, the compromise of the client to subvert the signing process and sign malicious artifacts; and third, the compromise of the certificate authority to conceal a malicious client.
Scope Provider Compromise (Types I & II)
If a scope provider were to be compromised (say, by hacking an OIDC provider), they can issue fraudulent scopes to impersonate a legitimate signer.
Under our system assumption, even if an attacker fraudulently obtains one valid identity token (say, by hacking an OIDC provider), they still lack the other required scopes – for example, they cannot fake the TEE-based attestation from the signing client. Because at least one of the multiple checks will fail, the attacker cannot produce a complete DiVerify proof to sign malicious code. Hence, a single compromised provider is not enough to break the chain of trust, thus providing security goal S1 and S2 guarantees
Client Compromise (Type III)
In this scenario, we identify the following cases:
  1. 1.
    Bypass authentication: A compromised client may skip signer authentication and spoof scopes, producing a DiVerify proof that enables unauthorized signing.
  2. 2.
    Substitution signing: Since the client controls the ephemeral signing key, it can misuse it to sign other software, yielding valid but malicious signatures.
Naively trusting such signatures allows attackers to distribute compromised software.
DiVerify’s resilience depends on the operating mode. In Legacy-Compatible mode, the client is trusted to authenticate signers, but without a TEE its behavior cannot be checked. A compromised client can spoof or omit verification steps, leaving this mode vulnerable.
In contrast, Core mode and Trusted Authentication mode delegate verification to a daemon running inside a TEE, while the untrusted client only forwards requests. Any modification of the daemon or its execution environment is detectable through attestation. Core mode validates the signing context at verification time, while Trusted Authentication mode validates it at certificate issuance. In both cases, signatures produced by a compromised client are detectable and rejected, thereby satisfying Security Goal S3.
A remaining risk is payload substitution: e.g., Alice asks to sign package foofoo, but the client forwards $foo^{\prime}$ to the daemon. Because authentication is decoupled from content, Alice has no visibility into what is actually signed. This can be mitigated if (i) the daemon displays the artifact information such as the file name and its cryptographic hash directly to Alice through a trusted channel (e.g., using GPU passthrough), without relying on the compromised client or (ii) a scope provider attests to the client’s integrity using TPM ME [32] or IMA [48].
Certificate Authority Compromise (Type III)
When a CA is used, and is compromised, they can issue signing certificates without validating the client’s attestation, or may deliberately fail to flag invalid attestation. This allows a rogue client signing to appear legitimate to downstream verifiers. As a result, malicious software may be signed and distributed with a valid certificate. DiVerify mitigates this risk to varying degrees depending on the operating mode.
Trusted Authentication Mode assumes CA trust but includes the client’s attestation within the certificate. If a verifier re-validates the embedded attestation independently rather than trusting the certificate blindly, it can detect that the CA failed to perform required checks holding CAs accountable. This attack is not applicable in Core Mode, as the CA is removed from the trust chain entirely. Instead, the verifier directly validates the client’s attestation before accepting a signature, eliminating the risk of CA compromise. Legacy-Compatible Mode offers no protection against this attack. Since attestations are not part of the signing process, there is no mechanism for verifying whether the CA observed valid scopes or verified the client’s integrity before issuing the certificate.
7.2 Mitigating Past Attacks with DiVerify
We illustrate how DiVerify could have mitigated the attacks earlier described in section˜3.2.
  1. 1.
    ESLint (Type I): The attackers used stolen maintainer credentials to publish a malicious package. With DiVerify, identity verification would have required additional contextual scopes, not just possession of the maintainer’s token. Since the attacker could not produce these scopes, their attempt to propagate the modified package via npm would have failed
  2. 2.
    Midnight Blizzard (Type II): The attackers leveraged OAuth tokens obtained through a compromised identity provider. If the relying services had required DiVerify’s contextual scopes beyond simply accepting valid OAuth tokens, the attacker would have been unable to authenticate through the necessary scope providers, and the attack would have been prevented.
  3. 3.
    Codecov (Type III): The tool was silently modified, compromising CI environments of users who fetched it. With DiVerify, the package maintainers would have published a policy defining the tool’s known trusted state. At runtime, the tool would generate an attestation, enabling verification against this trusted state. Since the modified tool would have deviated, users could have detected and flagged the compromise.
8 Implementation and Evaluation
In this section, we assess DiVerify ’s practicality by measuring signing/verification times and storage overhead. Given that our implementation is built on existing signing tooling, the only costs for adoption are those of storing and executing new processes and tools within an ecosystem.
8.1 Implementation
Table 2: Implemented Scope Providers(see section˜5.3 for type) Location indicates where provider runs. Size is lines of code via sloc with default parameters.
Provider Type Location Size
OIDC [11] VP Remote 46
Device Fingerprint SP Local 22
Security Key [90] SP Local 28
Table 3: Average Signing & Verification Overhead Introduced by DiVerify (in milliseconds). L1–L3 indicate trust levels determining the signer scopes required during signing.
Legacy Compat. (ms) Trusted Auth. (ms) Core (ms)
L1 L2 L3 L1 L2 L3 L1 L2 L3
Signing Signing Time 50.93 81.53 76.59 329.66 328.00 323.79 73.73 68.59 86.98
Quote Gen. 22.86 26.00 25.60 25.27 24.40 25.60
Fulcio 38.40 39.15 32.99 253.06 254.48 249.06
Verification Verify Time 1483.53 1357.36 1403.96 1336.44 1397.01 1381.50 1668.88 1604.10 1659.33
Quote Verif. 216.01 187.04 211.34
TUF 1287.25 1162.46 1230.16 1164.88 1217.34 1184.39 1287.41 1274.11 1295.11
We extended the securesystemslib interface for signing and verifying signatures [45] to prototype DiVerify capabilities by making four major changes:
First, we implemented a DiVerify-aware signing to support both signing client and verifier. The client is responsible for retrieving signer scopes, signing and generating remote attestation (section˜5.2.2, and section˜5.2.3); on the verifier side, it enforces scope checks against embedded proofs and package policy (section˜5.2.5). We also implemented separate daemon logic to support operational modes where scope retrieval and signing are delegated to a dedicated background service. We deployed a local RSTUF[72] to securely distribute DiVerify policy(section˜5.4). Second, we adapted the client to support the models described in section˜6. We adapted this client along other identity-based software signing tools such as sigstore-python [82] with similar arguments and parameters to ensure user familiarity.
Third, we utilized existing authentication protocols to provide user scopes. We use the token beacon [11] to request identity tokens from a GitHub OIDC provider. We implemented two static scope providers: a custom one for measuring device fingerprint, and another for attesting[90] the signer’s security key. See Table 2. We provide three bespoke configurations using these scope providers as tighter security settings (L1 to L3).
Finally, we customized the Fulcio certificate authority [81] to incorporate DiVerify proof verification prior to certificate issuance. This required modifying 201 lines of code. Verified proofs are embedded in the issued certificate.
To support remote attestation, we deployed a local Provisioning Certificate Caching Service [15], which supplies quote generation and verification services with Intel SGX collateral: Provisioning Certification Key certificates, Certificate Revocation List, Trusted Computing Base info, and Quoting Enclave identity data. For quote verification, we use Intel’s SGX DCAP Quote Verification tool [16].
Experiment Setup.
We ran our experiments using Docker containers on a 4.0 GHz Intel Xeon Platinum 8580 machine with 499GB of RAM and SGX support, reflecting a realistic industry-grade setup for scalable signing tasks [37]. TEE components were run using Gramine [12], which automatically handles the loading and execution of apps within the enclave.
8.2 Performance Evaluation
We evaluated DiVerify’s performance across the three modes, legacy-compatible, Trusted Authentication, and Core. All experiments used the sample trust level configuration shown in Listing LABEL:lst:diverify-config-structure.
We evaluated different policies to mode mapping with constraints to mitigate the attack vectors in our threat model. To evaluate DiVerify’s performance overhead, we averaged over 10 iterations of artifact signing and verification for each combination of mode/trust level.
8.2.1 Performance Overhead
Table˜3 presents breakdown of time spent across the signing and verification phases. The primary contributor to signing latency is the Fulcio CA. This is evident in Trusted Authentication Mode, where the CA verifies client integrity attestation leading to longer signing times compared to the Legacy, which skips attestation.
Signing using Core offers stronger security guarantees with minimal overhead. It can reduce signing time by up to 13ms compared to the legacy-compatible variant because all operations are handled within the daemon, avoiding round trips to an external CA. Still, both signing modes complete in under 100ms. On the other hand, verification in Core is more expensive, due to additional attestation verification performed by the verifier, resulting to about 11% average latency increase compared to the legacy-compatible mode.
Despite the added security guarantee, DiVerify’s performance remains comparable to state-of-the-art benchmarks [62]. On the verification side, the main cost stems from updating TUF metadata, which as specified, requires multiple round-trips to a remote repository. Core incurs slightly higher overhead than the other modes due to the additional verifier-side effort in validating the client’s integrity attestation. Nevertheless, the overhead remains within acceptable bounds and on par with current signing and verification tools.
8.2.2 Storage Overhead
We measured the size of DiVerify proofs across different operational modes, ranging from 0.4 KB (minimal identity attributes) to 6.7 KB (with full scope and attestation). This size remains fixed for deployments with consistent scope requirements. To understand the practical impact, we analyzed the top 20 most downloaded PyPI packages over the past month [70] and found an average package size to be 1.68 MB. From our sample, DiVerify adds at most 0.39% overhead, which is minimal relative to the package size. While the relative impact decreases for larger packages, the added size remains modest even for smaller ones.
9 Discussion
9.1 DiVerify & Signer Privacy
Our work thus far has focused on security and deployment objectives, not on anonymity or privacy for the signer. DiVerify, as described, ties a lot of identifiable information into the signature bundle, which might concern privacy-conscious signers or organizations. For instance, the certificate will reveal exactly which identity, devices, and scope providers were used. To address this privacy concern, DiVerify can integrate Speranza[52]. Speranza relies on a chain-of-custody proof (i.e., identity co-commitments) from a repository and an IdP. Integrating DiVerify with Speranza would require to adapt scope providers to provide a portion of an identity CoCommitment.
Consequently, while the set of providers remains visible, it becomes impossible to decipher the user who authenticated with them. This effectively provides privacy for signers, ensuring their anonymity when signing.
9.2 Applicability/Deployment Considerations
DiVerify’s design emphasizes backward compatibility (deployment goal (D1)). It builds on standard authentication flows and adds contextual identity “scopes” to signatures, allowing it to plug into existing code signing systems without replacing any infrastructure. Because of this, organizations can introduce DiVerify alongside current workflows; existing verification processes still work (even if they ignore the extra scope data). This means teams can roll out DiVerify incrementally – for example, starting with basic identity checks and gradually adding more stringent scope requirements as they become ready.
Integrating DiVerify requires little changes to existing tooling (satisfying D2). Our demonstration on secure systems library [45] illustrates this. The integration required only 2007 lines of code for the client side signing and verification tooling, and 201 lines of code increase for Fulcio. These modest changes implies maintainers can add DiVerify with very little development effort, keeping the codebase easy to manage.
Because DiVerify leverages existing capabilities, maintainers don’t have to reinvent their environment, just augment it. Another consideration is the choice of scopes for a given context. Our approach is flexible across the claim types identified in 2.2. In automated build systems, DiVerify could be used to ensure that build and signing servers themselves authenticate via scopes (like machine certificates or runner attestations), and that triggers for signing (like a GitHub Actions workflow run) carry a scope proving an authorized user initiated them.
10 Related Work
This work relies on existing works from access control mechanisms, as well as distributed identity mechanisms. We contextualize DiVerify in these techniques, focusing on their applications to software supply chains.
Access Control Mechanisms
Role-Based Access Control (RBAC) has long been foundational in managing permissions [77], with widespread adoption in general-purpose systems [57, 56], cloud platforms [41, 2], and social coding environments [21, 22]. In supply chain contexts, Kuppusami et al. [43, 23] propose repository-specific models for permission management, leveraging abstractions like branches and commits. DiVerify differs by mapping permissions across heterogeneous domains without sacrificing security guarantees, particularly in the identity-based software signing context.
Several studies have also identified limitations in RBAC implementations for supply chain security [58, 39], focusing on repository metadata and CI platforms. These insights inform DiVerify’s design, especially in modeling semantics across diverse systems.
Verifiable Identities and Credentials
Similarly, the application of identity for different contexts has surfaced in general identity work. Perhaps one of the richest lines of work in this regard is that of Verifiable Credentials. Work from [34, 50, 59, 47, 66] introduces the ability to disclose, verifiably, that a property is held by a credential holder. DiVerify provides a similar goal by building upon existing and widely deployed technologies (such as OpenID Connect). Properties such as the ability to commit to a repository or holding ownership of a project can be achieved by off-the-shelf algorithms and tools.
Similarly, applications of distributed identity frameworks (i.e., DiD [89, 5, 88]) can be used in lieu of identity providers. This has been considered in standards such as SCITT as a possible building block for a global software supply chain ledger [29]. Were these standards to become widespread, DiVerify can leverage them. However, it is still necessary to apply the policy logic to merge various providers described in this paper (section˜5.3) to ensure proper capabilities are applied.
Threshold-based Authentication
Threshold-based authentication schemes, which distribute trust across multiple identity providers [33] using secret sharing [80], have been proposed for enhancing resilience and privacy in federated identity systems. These approaches align with DiVerify’s goal of federated identity verification but typically rely on a trusted dealer. DiVerify avoids this by ensuring no single party can subvert trust, maintaining a decentralized model
11 Conclusion
We presented DiVerify, a novel approach to strengthening software signing systems against threats from both identity provider and client compromise, two key problems in current signing architectures. DiVerify addresses these by collecting diverse, verifiable trust signals for both signers and clients, enabling quick detection of malicious signing and improving the resilience of software signing as a security measure.
Our prototype implementation demonstrates that DiVerify is both practical and effective, enabling proactive threat mitigation with minimal performance overhead. DiVerify establishes a security baseline that signing tools can adopt to provide robust code-signing guarantees. We leave integration with build systems and measuring usability as future work.
Appendix
The appendices contain the following material:
  • Appendix˜A: An extended version of the assumptions we made.
  • Appendix˜B: Sample DiVerify Artifact Structure
  • Appendix˜C: Sample policy used in our analysis, with signer scope mapped to the expected claims, and the constraint the signature must meet to be trusted.
Appendix A Extended Assumptions
Beyond the attacker models described in section˜3.2, we make standard assumptions from previous literature [7, 62]. First, attackers are able to interpose communications, but standard mechanisms to prevent person-in-the-middle attacks (e.g., TLS) are in place and all parties utilize these. Similarly, attackers are able to replay communication messages to attempt to impersonate roles in the system.
However, an attacker is not able to:
  1. 1.
    Break standard cryptographic algorithms (e.g., sha256, ed25519). An adversary capable of breaking standard cryptographic algorithms is out of scope.
  2. 2.
    Overcome kernel-level or TEE-based system measurement and isolation techniques. While attacking TEEs is a thriving area of offensive security research, we assume that these systems are functioning correctly.
  3. 3.
    Compromise a collection of IdPs. While a sophisticated attacker could compromise large swaths of identity infrastructure, this is rarely the case in practice due to the practical difficulty of breaching multiple independent scope providers at once, especially when operated by different organizations with diverse infrastructures and security postures.
Lastly, we assume that the Signer is legitimate (i.e., behaves honestly and does not intend to signing of untrusted code). We also assume that the verifier behaves honestly and correctly verifies signatures and associated identity information according to the expected protocol.
Appendix B Other DiVerify Artifact Samples
Listing LABEL:lst:diverify-config-structure defines a sample trust level configuration used in DiVerify, specifying the scopes required for a signer at each level. Level 1 provides identity claims, level 2 adds possession claims, and level 3 includes contextual claims in addition to level 2.
Listing LABEL:lst:proof-structure shows a sample structure of the DiVerify proof, showing the sample collected scopes that represent claims about both the signer and the client.
1{
2 "Level": 1,
3 "identity": {
4 "oidc": true
5 },
6 "Level": 2,
7 "identity": {
8 "oidc": true,
9 "security_key": true
10 },
11 "Level": 3,
12 "identity": {
13 "oidc": true,
14 "security_key": true,
15 "device_fingerprint": true
16 }
17}
Listing 2: Sample Trust Level
1{
2"config_version": "1.2",
3"trust_level": 3,
4"identity": {
5"oidc": {
6 "identity": "...",
7 "provider": "...",
8 "token_hash": "...",
9 },
10"device_fingerprint": "...",
11"security_key": {
12 "slot9a_public_key": "...",
13 "slotf9_attestation_cert": "..."
14 },
15 },
16"signing_key": "..."
17"remote_attestation": {
18 "enclave_quote": "..."
19 }
20}
Listing 3: Sample DiVerify Proof
Appendix C Test DiVerify Policies
Figure˜5 presents a sample policy used to evaluate whether to trust a package signature.
Figure 5: An example DiVerify policy specifying expected claims. The final field is a Boolean rule constraint under which a package signature is to be trusted. This policy is defined by the software producer and enforced by the verifier, ensuring that only signatures that meet this constraint are accepted, which strengthens the overall trust in the signing process.