### v3

# DiVerify: Hardening Identity-Based Software Signing with Diverse-Context Scopes

Abstract.
Identity-based code signing enables software developers to digitally sign their code using cryptographic keys. This key is then linked to an identity (e.g., through an identity provider), allowing signers to verify both the code’s origin and integrity. However, this code–identity binding is only as trustworthy as the mechanisms enforcing it. State-of-the-art identity-based code signing schemes present a major shortcoming: these schemes fail to provide verifiable information about the context in which a signature is generated. This verifiability is crucial given that modern attackers have subverted long-established security assumptions, namely, that the identity provider ecosystem, as well as signing software itself is trusted.
To address these issues, this paper introduces a diverse identity verification framework, DiVerify, that distributes identity-based verification across multiple entities and enforces stronger guarantees about the signing context. DiVerify makes it possible to provide end-to-end verifiability of not only a signer’s identity (via multiple such signals), but also a signer’s software stack (e.g., to verify no malware is present on a system at the time of signing). DiVerify is aimed at deployability, and leverages a meta-protocol to gather various trust signals and a binding mechanism to address the aforementioned, novel software supply chain attack vectors. We evaluate DiVerify’s performance and confirm it is cheap to deploy and non-intrusive to developers: it only incurs a few kilobytes of additional storage (less than 0.4% of the average package size in widely used ecosystems like PyPI), and signing completes in under 100ms on a server-grade deployment.
“Yo soy yo y mi circunstancia.”
“I am I and my circumstance.”
— José Ortega y Gasset
Software Signing, Software Supply Chain Security, Policy Transparency
$c$s: Security and privacy Trust frameworksccs: Security and privacy Software security engineeringccs: Security and privacy Distributed systems security
1. Introduction
Software supply chain attacks are a major, widespread concern due to their large-scale impact on software users’ safety (Reuters, 2021; Rapid7, 2021; CNCF TAG-Security, 0 12). These attacks involve the injection of malicious code into a software artifact and subsequent exploitation in downstream systems (Okafor et al., 2022). In response, government, industry and academia have proposed means to improve provenance along the software supply chain (Cooper et al., 2018a; House, 2021; Foundation, 2021).
The most common provenance mechanism for in this setting is software signing (Rivest et al., 1978; Kalu and Davis, 2025). Software providers attach a digital signature to an artifact using public key cryptography, and software consumers can then test this signature to detect tampering (Cooper et al., 2018b). Despite decades of use (Rivest et al., 1978; Diffie and Hellman, 2022), adoption has been limited by usability (Whitten and Tygar, 1999; Kalu et al., 2025a) and key management challenges (Schorlemmer et al., 2024; Kalu et al., 2025b). To address this, identity-based signing platforms have emerged (Schorlemmer et al., 2025), still grounded in public key cryptography (Newman et al., 2022; Lorenc, 2023), but linking signatures to verified identities via identity providers (IdPs) and short-lived ephemeral keys to allow a single signing event.
However, the state-of-the-art identity-based signing solutions are not resilient to compromises within their underlying trust model. If any party involved in the signing operation is compromised, then the provenance guarantees are lost. This limitation manifests in two ways. First, identity-based signing systems rely on a single IdP to identify a signer and issue authentication tokens used for signing – if this IdP is compromised, it renders all associated signing events untrustworthy. Many recent attacks have targeted identity providers, including the Midnight Blizzard intrusion (Microsoft, 2024), Apple’s SSO (Jain, 2020), and attacks on Facebook (Poza, 2018; Rosen, 2018). Second, signing clients responsible for brokering interactions between the software author, the identity provider, and the repository hold broad authority during the validity period of their ephemeral keys. If compromised, they can be coerced into producing valid signatures over malicious software. The Diamond Sleet (ZINC) attack (Microsoft, 2023), which abused code-signing system’s privileges to sign malicious variants of an application, exemplifies this risk. Current systems make an undue assumption that both Identity providers and clients will never be compromised.
In order to surmount this limitation, identity-based signing must provide verifiable evidence regarding the context in which signing took place. Verifiers should confirm how the signer authenticated or whether the signing client is not compromised. This capability addresses both problems outlined above. In sum, by forwarding verifiable evidence about the security posture of the signing process, verifiers can ensure that the signer is trusted, and that the signing event was carried out in a way that can be trusted as well.
Although many ad-hoc mechanisms exist to harden identity verification such as multi-factor authentication (Golla et al., 2021; Reese et al., 2019), hardware-backed credentials (Yubico, [n.d.]; Alliance, 2022), device-binding schemes (Hodges et al., 2021), and even cross-device secret-sharing approaches (Laing et al., 2022), state-of-the-art identity-based signing systems do not provide verifiers with verifiable evidence of these conditions. Similarly, client integrity can be reinforced through attestation (Kim et al., 2015) or other local hardening techniques (Wilkins and Richardson, 2013), these assurances never reach the verifier. And without evidence, the verifier cannot distinguish a legitimate signing activity from one carried out on a compromised client, leaving attackers free to exploit valid credentials under malicious conditions. We posit that this ability to transparently transfer this information to the verifier is crucial to addressing supply chain attacks targeting trust infrastructure (Laurie, 2014).
In this paper we present DiVerify, a framework for identity-based signing that addresses these challenges, and a prototype system that instantiates it. DiVerify extends existing system threat model and introduces two main components. The first is the DiVerify proof, a cryptographically verifiable bundle produced for each signing event that binds together identity attributes, authentication context, and attested measurements of the signing environment. Our construction provides a meta-protocol that integrates common, reliable mechanisms. Our meta-protocol combines these heterogeneous signals into a homogeneous proof that exposes independently-verifiable trust assertions. A second component, a DiVerify policy, specifies the conditions under which a verifier should accept a proof — e.g., dictating the minimum authentication strength, acceptable identity claims, and expected trusted execution environment integrity — and enables these conditions to be checked independently of the upstream signing pipeline. These policy constructions are essential to accommodate developer signing practices and provide expressions that can integrate verifiability for these heterogeneous signals. Together, these allow verifiers to enforce explicit, context-aware trust requirements rather than inheriting implicit assumptions identity providers or signing clients make.
The design of Diverify accounts for providing verifiability even as trust expectations change. Developers replace devices, add authentication factors, or migrate across platforms, and maintainers may strengthen or revoke scopes, making communication shift in DiVerify policies is paramount. Thus, DiVerify introduces a principled mechanism for representing and evolving verifier trust in a controlled and verifiable manner, ensuring that trust decisions remain consistent even in an adversarial environment.
We evaluate DiVerify against the adversary in our threat model and show that DiVerify strengthens the security posture of the systems by increasing the adversarial effort required for compromise. We also examine how different DiVerify instantiations balance security and performance trade-offs when integrated into existing signing systems. To demonstrate practical relevance, we analyze several historical supply-chain attacks and show that DiVerify could have detected or prevented them. Finally, we measure the performance overhead of our prototype and find that signing completes in under 100 ms and verification adds only 11% latency compared to legacy-compatible approaches, highlighting its feasibility for real-world adoption.
In summary, we make the following contributions:
  1. (1)
    Based on recent attacks, we extend the existing identity-based signing system threat model. We then propose a novel verifier-centric trust model for identity-based code signing. We treat signing context, identity, authentication, and signing-environment integrity as a single, verifiable proof.
  2. (2)
    We develop mechanisms for authenticated, evolvable policy state that allow verifiers to detect replayed or downgraded trust configurations and reason about policy-level attacks not covered by existing identity-based signing systems.
  3. (3)
    We design and implement DiVerify, a system that enforces context-aware signing policies over these proofs, replacing ad hoc verifier configuration with an explicit policy model.
  4. (4)
    We evaluate DiVerify’s security and performance, measuring the cost of generating and verifying proofs and showing that DiVerify can mitigate several real-world attacks that prevailing solutions could not.
Significance
Identity-based software signing has become a central security primitive across modern software ecosystems, providing the attribution signal upon which downstream trust decisions depend. This model underlies workflows in cloud platforms (The Kubernetes Authors, 2 11), AI and MLOps model pipelines (Sablotny, 2025; Font, 2025), and package ecosystems (npm, Inc., 5 22; Ingram, 2024). DiVerify reinforces this primitive by reducing reliance on any single authority and by mitigating compromise of the signing primitive itself. Although we instantiate DiVerify for software signing, the underlying model generalizes to any system where provable attribution is the foundational security property upon which all downstream trust relationships depend.
2. Background
To describe DiVerify, we first discuss its intended role: code signing in the software supply chain (section 2.1). The DiVerify protocol also requires understanding authentication mechanisms (section 2.2) used to establish identity and context, and remote attestation, which we leverage in our implementation.
2.1. Software Supply Chain and Code Signing
The software supply chain involves all the activities carried out in the task of creating and distributing software, such as coding, compiling, testing, packaging and deployment (Schorlemmer et al., 2024; Okafor et al., 2022). These activities rely on a wide range of tools, libraries, frameworks, and services, creating a complex web of dependencies (Kumar et al., 2017; Synopsys, 2023a; Kumar et al., 2017; Synopsys, 2023a). This complexity makes it a prime target for attackers, as seen in incidents like SolarWinds (Google Cloud, 2020) and Log4j (National Cyber Security Centre, 2021). As a result, organizations are increasingly adopting practices like code signing, improved dependency management and vetting (Williams et al., 2025; Pashchenko et al., 2020; de Souza and Redmiles, 2008), continuous monitoring (Synopsys, 2023b), zero-trust architectures (Lamb and Zacchiroli, 2022; Torres-Arias et al., 2019; Amusuo et al., 2025), and Software Bills of Materials (SBOM) (National Telecommunications and Information Administration, 2021; Abu Ishgair et al., 2025) to ensure transparency, integrity, and compartmentalization throughout the supply chain.
Code signing allows a verifier to establish trust with a software producer (Rivest et al., 1978; Kalu and Davis, 2025). The process uses a keypair $(k_{p},k_{s})$ and an artifact $(SW_{a})$ to generate a signature $(S_{sw})$. A verifier will then use $(K_{p},SW_{a},S_{sw})$ to verify whether $(SW_{a})$ is, first, integral (i.e., hasn’t been tampered with), and second, belongs to a trusted party. Traditional tools like OpenPGP (OpenPGP Project, [n.d.]) use manual key management and user-controlled trust models (Chandramouli et al., 2013; Unger et al., 2015), requiring developers to generate, safeguard, and distribute long-lived keys, introducing usability challenges and risks of compromise. Identity-based solutions such as Sigstore (Newman et al., 2022), OpenPubKey (Heilman et al., 2023), and Keyfactor (Keyfactor Team, 2025) simplify this by using ephemeral keys and identity federation protocols (e.g., OIDC, OAuth, SAML). These allow developers to prove identity ownership via standard authentication flows, then linking identities to signing keys through certificates or embedded tokens.
2.2. Authentication Mechanisms
Identity-based signing uses authentication to verify the identity of a signer. Authentication mechanisms vary, but generally fall into three categories (Huang et al., 2010; Mahadi et al., 2018): identity, possession, or context.
Identity claims verify who a signer is, using protocols like OAuth 2.0 (OAuth.net, 2 05) and OpenID Connect (OpenID Foundation, 2023). These protocols issue signed tokens (e.g., JWTs (Jones et al., 2015)) from trusted providers like Google. Distributed Identities (DiD) (World Wide Web Consortium, 2022) and CryptID (Cryptid, [n.d.]) provide mechanisms similar to OIDC but act as a meta-protocol for identity-specific claims.
Possession claims prove control over physical elements such as a hardware token, smartcard, or cryptographic key. The authenticated party demonstrates possession of a secret (e.g., a private key) through a cryptographic challenge-response protocol. Protocols like FIDO2/WebAuthn, use devices (e.g., YubiKeys (Yubico, [n.d.])) to protect the private key and perform challenge responses locally ensuring that only someone with the device can authenticate.
Contextual claims assert properties of the execution environment, focusing on system state rather than identity or key possession. They support trust decisions using runtime conditions like device or software integrity. Techniques include device fingerprinting (Rizzo et al., 2021), TPM-based measurements (Microsoft, 4 21; Intel Corporation, 2025), and kernel-level attestation via Linux’s Integrity Measurement Architecture (IMA) (Linux Integrity Subsystem, 4 21).
DiVerify extends identity-based signing by supporting possession and contextual claims in addition to identity claims, allowing verifiers to incorporate evidence about the signing environment and execution context into acceptance decisions.
2.3. Trusted Execution Environments
Remote Attestation enables verifiers (e.g., a remote user or system) to confirm software execution within a secure and trusted environment on a remote machine. Trusted Execution Environments (TEEs) enable this capability by isolating processes using hardware mechanisms like secure memory and privileged execution contexts (Arm Ltd., 4 21).
This is achieved through a cryptographic protocol in which the remote system generates a signed attestation report or quote that includes a measurement (e.g., hash) of the software and configuration loaded into a secure execution environment. The verifier can compare this measurement against a known good value and determine if the software executing is indeed the expected one.
Trust in remote attestation relies on hardware-rooted chains of trust, such as those provided by Intel Software Guard Extensions (SGX) (Costan and Devadas, 2016), or Trusted Domain Extensions (TDX) (Intel Corporation, 2021) which include embedded attestation keys and associated endorsement certificates issued by the hardware manufacturer. These cryptographic materials ensure that only legitimate, manufacturer-approved hardware can produce valid attestation signatures.
3. Parties, Roles, and Threat Model
We formalize identity-based software signing systems and the context in which they operate (section 3.1). We show that existing threat models are too weak relative to recent attacker behavior and define the security goals of both identity-based software signing in general, and DiVerify in particular (section 3.2).
3.1. Parties and Roles
We consider a standard (Grassi et al., 2017) identity-based software signing setting with the following parties interacting with each other:
Signer: an entity (i.e., a developer or an automated system) who intends to generate a signature $(S_{s}w)$ over an artifact.
Client: a tool or service with which the Signer interacts to produce such signatures. To do so, a Client needs to know the Signer’s identity to complete a signing request. This request involves generating key material and interacting with an Identity Provider to establish trust in such material.
Identity Provider (IdP): an entity that authenticates the Signer through the Client and provides qualified attributes about the User. In DiVerify we rely on a general-purpose instantiation of an IdP we call scope providers as described in section 5.3.
Package registry: a storage service for software packages and their signatures. Users publish, share, and retrieve artifacts for distribution and verification, e.g., NPM (JS) or Hugging Face (ML).
Verifier: is the counterpart of the Signer. It ensures that the signed software is valid by validating a signature against a software artifact using a certificate or public key.
Figure 1. State-of-the-art identity-based code signing system model, highlighting the types of attacker capabilities that such systems are not resilient against. These attack types are discussed in section 3.2.
A typical high-level interaction flow is illustrated in Figure 1. There are five steps. (1) The signer initiates a signing request by interacting with a client. (2) The client redirects her to an IdP to authenticate. (3) The signer authenticates, an identity token is obtained, and sent to the client. (4) Optionally, the client requests a signing certificate (which binds the identity to a signing key) from a certificate authority (or CA). (5) Finally, the client signs the software and uploads the signature to a registry.
A successful flow allows a verifier to obtain a $(S_{sw},K_{p},SW_{a})$ tuple from a source (e.g., a registry) and ensure the artifact is not tampered with. The verifier must also evaluate the identity attributes associated with the public key $K_{p}$ such as the signer’s email address and the issuing identity provider against locally configured trust expectations. These identity checks determine whether the signature corresponds to an authorized signer and thus whether the artifact should be accepted.
3.2. Threat Model
We consider a stronger threat model than previous literature to reflect modern attacker behavior. Components traditionally treated as trusted are now realistic targets whose compromise directly undermines security guarantees.
Prior work on the software supply chain has modeled threats to specific segments of the chain, such as source-code integrity (Vu et al., 2021; Gonzalez et al., 2021), build-system compromise (Fourné et al., 2023; Ohm et al., 2020), maintainer account compromise (Zimmermann et al., 2019; Samuel et al., 2010; Gu et al., 2023; Huang et al., 2024), package registry (Kuppusamy et al., 2016; Zimmermann et al., 2019; Zahan et al., 2022; Wu et al., 2023; Nikitin et al., 2017), mirror compromise (Cappos et al., 2008), package-distribution (Gao et al., 2024) and dependency-resolution attacks (Neupane et al., 2023; Vu et al., 2020). These models assume a trusted identity provider and signing client and rely on log transparency (Al-Bassam and Meiklejohn, 2018; Blauzvern, 2023; Laurie, 2014) to detect misbehavior after an event has been maliciously signed.
For instance, Log Transparency, places the responsibility for detecting misuse on signers, while the security consequences fall on downstream users. Because detection is retrospective, any delay between compromise and discovery exposes verifiers to the risk of accepting malicious artifacts. Yet verifiers must make trust decisions at verification time, and current systems provide them with no direct evidence about whether a signature reflects legitimate authentication or signer intent.
Recent incidents have shown that these components are realistic targets for attackers (Microsoft, 2024; Jain, 2020; Poza, 2018; Rosen, 2018). Excluding these capabilities yields a threat model that is too weak to capture the risks faced by modern software supply chains.
This motivates a threat model in which verifiers can validate these conditions themselves, rather than depend solely on retrospective mechanisms that upstream parties may not notice or act upon. We therefore extend prior threat models to treat the identity and signing ecosystem as an attack surface, capturing adversaries who can produce valid signatures without the legitimate signer’s intent. Regardless of their capabilities, all threat models in this domain assume that attackers aim to cause verifiers to accept tampered software. To do so, they may impersonate, tamper with, and/or attack any of the roles in the signing workflow, or manipulate the verifier’s locally configured trust expectations.
Attacker Goal
Ultimately, for all these threat models, attackers are successful if they can cause a verifier to accept a malicious artifact $(SW_{a})$. We want to raise the cost for such compromises by increasing the evidence an attacker must produce to induce acceptance.
Attacker Capabilities
We describe three main types of adversaries with the following capabilities, reflecting the different ways an adversary can obtain signing authority:
  1. (1)
    Type I: Credential Compromise. The attacker compromises a signer’s account, typically through credential theft (e.g., phishing, password leaks, session hijack). Once the attacker has access to the account, they can use it to authenticate and sign software on behalf of the signer without their consent. Type I attackers are successful if they are able to identify themselves on behalf of the Signer against an IdP.
  2. (2)
    Type II: Compromised or Rogue Identity Provider. The IdP is compromised or malicious, and can issue identity tokens for a signer who did not authenticate. An attacker can then use the fraudulently issued token to impersonate the signer gaining signing access. Type II attackers are able to impersonate an IdP and thus a Signer without interaction with either.
  3. (3)
    Type III: Compromised Signing Tool. The attacker can tamper with the client software to sign software beyond what the user originally intended, allowing unauthorized signatures.
These attacks are not theoretical, as shown by the events below:
  1. (1)
    ESLint (Type I: Credential Compromise): ESLint, a widely used JavaScript linter, was compromised in 2018 when attackers used stolen npm credentials to upload malicious package (Zhu, 2018). Because the system relied only on token-based authentication, it could not tell legitimate actions from malicious ones.
  2. (2)
    Midnight Blizzard (Type II: IdP Compromise): In 2023, the nation-state actor Midnight Blizzard compromised a legacy Microsoft test account to create malicious OAuth apps with elevated privileges (Microsoft, 2024). These apps granted persistent access to Exchange Online, bypassing identity checks. Since the tokens were issued by a trusted provider, the system lacked the context to detect the compromise.
  3. (3)
    Codecov (Type III: Tool Compromise): A widely used code coverage tool integrated into CI was compromised in 2021, when attackers added code to exfiltrate sensitive information (Codecov, 2021). The attack persisted undetected as users implicitly trusted the tool, with no way to verify its runtime state. Other attacks have also taken this form of abuse (Thomas et al., 2016; Kotzias et al., 2015; Kwon et al., 2017).
Trust Assumptions
We assume legitimate signers do not sign untrusted code, verifiers execute verification procedure correctly, the trusted execution environment setup remains secure, and not all Scope providers can be compromised at the same time. Rationales and standard assumptions are discussed in Appendix A.
4. System Goals and Overview
We now describe our deployment and security goals and provides an overview of DiVerify’s signing and verification flow.
4.1. System Goals
In order to address the attackers described in our Threat Model, DiVerify extends authentication beyond traditional credentials and hardens the security posture of signing clients; and limits reliance on any single identity provider by requiring multiple, independent sources of verification. We formalize these notions into these security goals:
  1. S1
    Resilience Against Identity Theft: Signing events are allowed only after verification from multiple independent sources, which limits trust assumptions on individual identity providers. This way, if an attacker compromises a Signer’s credentials, the attacker is unable to complete a signing request without a collection of these.
  2. S2
    Resilience Against IdP compromise: In addition to diversity of credentials, a Signer can request multiple identity providers to prove a diverse identity claim. This ensures that a compromised provider cannot impersonate its user unless it has also compromised a threshold number of other independent providers. Unlike S1, which protects against attacker control of a signer’s own credentials, S2 protects against an identity provider issuing fraudulent assertions even when the signer’s credentials are uncompromised.
  3. S3
    Resilience Against Compromised Clients: It should be possible to cryptographically verify that a signature was generated by a trusted tool and not a compromised version. This is possible by having signing information include information about the signing context as well as identity claims.
In addition to the security goals, DiVerify is designed with deployability in mind. We describe in section 10.2 how its design enables integration into current code-signing systems with minimal changes.
4.2. Components
Scope Providers
manage and issue claims about a signer’s identity, attributes, or environment context. Each operates within a distinct trust domain, that is, an independently managed authority responsible for issuing and validating its own set of claims about signers. They follow an ad-hoc protocol to verify properties such as control of an email address, possession of a cryptographic key, or ownership of a device and then issues a signed statement attesting to that property. DiVerify does not propose new protocols for scope providers, but builds on existing standards whenever possible. Scope providers must be able to authenticate themselves to the tool and prevent scope replay to ensure that claims are both trustworthy and fresh. When native protocols the required properties, DiVerify “wraps” them using a meta-protocol ( section 5.3).
Client
is a signer-facing component responsible for initiating signing requests. This can be instantiated as a command-line tool or library integrated into developer environments. The client operates within the signer’s local context and has access to the artifact to sign. It serves as the point of integration between the signer and the rest of the DiVerify architecture. Because the client runs in a potentially untrusted setting, we split the functionality of a traditional signing client into two parts: a front-end Client and a privileged Daemon.
DiVerify Daemon
is a long-running process responsible for coordinating with scope providers and performing signing operations. Importantly, it runs inside a TEE on the user’s machine. The daemon thus serves as a secure execution environment that can verify scopes and construct proofs with a higher level of assurance, being isolated from the rest of the system. It can also generate a remote attestation to prove its own integrity to external verifiers.
Verifier
is an end-user side component responsible for validating signed artifacts and assessing their trustworthiness. It operates independently of the signing process and has no privileged relationship with the signer.
4.3. System Overview
DiVerify introduces three mechanisms to strengthen identity-based software signing: (1)zero-trust signing architecture, provides verifiers with a DiVerify proof that binds diverse identity and client-trust conditions into a verifiable object (2) a DiVerify policy that gives verifiers explicit and evolvable control over the trust conditions under which signatures are accepted; Figure 2 illustrates how each mechanism integrates into the signing workflow.
These mechanisms are built on a disaggregated trust model that replaces traditional reliance on a single identity provider with a set of independent scope providers. Each scope provider asserts a specific property about the signer such as the identity, the device used, or the presence of another factor. When combined with attestation of the signing client’s integrity, these claims require an attacker to obtain strictly more and qualitatively different capabilities in order to cause a verifier to accept a malicious artifact, compared to signing based on a single identity assertion.
As shown in Figure 2, the signing process begins as in conventional identity-based software signing systems when requests that their client produce a signature over their software artifact. However, instead of obtaining a single identity token from an identity provider, the client quires multiple scope providers to assemble the set of claims required by the verifier’s trust expectations. e.g., a signer may obtain an OIDC token for identity, a fingerprint from the device being used, and a second factor assertion from a hardware token. In parallel, the client collects integrity evidence about its own execution environment.
The client aggregates these claims into a DiVerify proof. The proof captures a diverse security posture of the signing event and depending on the integration path (discussed in section 7), this proof is be distributed alongside the signature or embedded within existing certificate produced by identity-based software signing tools.
On the verifier side, trust decisions are guided by a DiVerify policy, which specifies the scopes, attestation properties, and the logical conditions that must hold for a signature to be trustworthy. e.g., a policy for Alice may require that software be signed by alice@software.sh, from a known device and using a registered hardware second factor. When Bob retrieves a signed artifact, he validates the signature and checks that its its accompanying DiVerify proof satisfies the policy. The policy therefore replaces the implicit trust delegation traditionally provided by identity providers with an explicit specification of the trusted signing conditions.
Figure 2. DiVerify overview. Scope providers assert signer claims that, together with client integrity attestation, form DiVerify proof. Verifiers check this proof against a policy to decide whether to accept the signed artifact.
As an example, suppose Alice maintains a software package and wants verifiers to trust updates only if she signed them as alice@software.sh, from her MacBook Pro, and using her registered Yubikey. She expresses this intent in a DiVerify policy. During signing, the client collects the required claims from the relevant scope providers and incorporates them into the DiVerify proof. When Bob later verifies the package, he checks that the signature is valid and that the accompanying DiVerify proof satisfies Alice’s policy. An attacker who merely obtains Alice’s password or steals her identity token cannot satisfy these additional contextual requirements, the proof fails the policy check and the signature is rejected. In this way, DiVerify prevents attackers from signing on Alice’s behalf even if some authentication factors are compromised.
We defer the discussion of policy distribution, evolution and revocation to section 6.2, which describes how verifiers obtain and ensure they use the right policy.
5. Generating and Verifying Proofs
DiVerify executes a two-phase protocol to sign and verify software. First it asserts and aggregates scope into proof describing security posture when signing was performed. Second, it validates the scopes and the artifact against some expected state.
5.1. Generating DiVerify Proofs
Proof generation proceeds as a single signing session in which the daemon (i) collects required scope claims, (ii) validates them, and (iii) binds them to the signing key and attestation.
DiVerify client advertises the set of scope providers they support. When the DiVerify daemon starts, it establishes communication channels with the client and with each supported scope provider.
The signing protocol, shown in Algorithm 1, begins when the signer initiates a signing request via the client, specifying the artifact identifier (e.g., a filename or container image tag). The client then forwards this request to the DiVerify daemon.
The daemon determines which scopes are required based on the supported providers. These scopes are independent, and may be obtained in parallel. The daemon generates a nonce to bind all subsequent scope responses to the current signing session. For each required scope provider, the daemon executes a scope request meta-protocol.
Algorithm 1 DiVerify Daemon Flow to Sign Payload
1:In: Artifact AA
2:Requires: Scope Provider $P$,
3:Out: ArtifactSignature $S$, DiVerify Proof dvpdvp
4:procedure SignRequest(AA)
5:  $scopes\leftarrow\{\}$
6:  for $p\in providers$ do
7:   $scopes[p]\leftarrow\text{AuthenticateSigner}(p)$
8:  end for
9:  if not ValidScope(scopesscopes) then
10:   return Err
11:  end if
12:  $sk,pk=\text{GenerateKeyPair}()$
13:  $S\leftarrow\text{Sign}_{sk}(A)$
14:  $custom\_data\leftarrow\text{Sign}_{sk}(hash(scopes))$
15:  $quote\leftarrow\text{GetQuote}(custom\_data)$
16:  $dvp\leftarrow\text{MakeDiVerifyProof}(scopes,quote,pk)$
17:  return $S$, dvpdvp
18:end procedure
For each scope request, the daemon sends a request of the form $request(n,claim)$ to the provider. The provider returns a tuple $\langle\mathit{sign(proof)},\ \mathit{proof}\rangle$, where $\mathit{proof}$ is a protocol-specific claim (e.g., an oidc token) that embeds the nonce, and the accompanying signature authenticates the response.
After collecting responses from all required providers, the daemon verifies each response by checking the provider signature and that the embedded nonce matches the current signing session.
If all scopes validate successfully, the daemon generates an ephemeral signing key pair and signs the artifact using the private key. The daemon then signs a hash of the validated scopes with the signing key and embeds the result into the custom data report of the quote produced by the TEE.
Finally, the daemon constructs a DiVerify proof consisting of the validated scopes, the signing public key, and the remote attestation. The daemon returns the artifact signature and the DiVerify proof to the client, which uploads them to the package registry. Embedding a signature over the validated scopes into the attestation quote binds the signing key and scope set to the TEE’s execution context, preventing replay of the attestation
Algorithm 2 Artifact Verification Flow
1:Input: Artifact A, Policy, ArtifactSignature $S$, DiVerify proof dvpdvp
2:Output: Boolean verification result
3:procedure VerifyArtifact(A, S)
4:  quotequote $\leftarrow$ $dvp.\text{quote}$
5:  custom_datacustom\_data $\leftarrow$ $quote.\text{custom\_data}$
6:  pkpk $\leftarrow$ $dvp.\text{pubkey}$
7:  scope_digestscope\_digest $\leftarrow$ $hash(dvp.scopes)$
8:  if not ValidQuote(quotequote) then
9:   return false
10:  end if
11:  if not VerifySignature(scope_digestscope\_digest,custom_datacustom\_data,pkpk) then
12:   return false
13:  end if
14:  if not EvaluatePolicy($dvp.scopes$, policypolicy) then
15:   return false
16:  end if
17:  if not VerifySignature(AA,$S$,pkpk) then
18:   return false
19:  end if
20:  return true
21:end procedure
5.2. Verifying DiVerify Proofs
To verify a signature, a verifier needs the signed artifact, the artifact signature, the DiVerify proof, and a verification policy specifying the expected signer claims and acceptance rules, defined in section 6.
The verification protocol follows Algorithm 2. The verifier first retrieves the policy. It then extracts the attestation quote, the embedded custom data, public key, and scope claims from the DiVerify proof. The verifier first verifies the integrity of the attestation quote and then checks that the scope digest matches the custom data using the extracted public key. Next, it evaluates the extracted scopes against the policy. Finally, it verifies the artifact signature using the public key. If any check fails, the verifier rejects the artifact.
5.3. Integrating Scope Providers
DiVerify requires that scope providers must be able to authenticate themselves to the daemon and prevent scope replay. Not all providers natively support these properties. To accommodate a wide range of provider capabilities, DiVerify supports two integration strategies; wrapping and TEE-based integration. We explore different provider categories and their are support in DiVerify.
Verifiable Providers (VP) can cryptographically authenticate themselves and issue scopes that are bound to a specific context, making them verifiable and replay resistant. A representative example is an OIDC token signed by a known Identity Provider such as Google. VPs are natively supported in DiVerify; integrating them requires only defining the appropriate syntax for policy verification.
Static Providers (SP) is not replay resistant, but the authenticity of the provider can still be verified through cryptographic signatures or secure communication channels. SPs are often stateless protocols that provide a fixed value whenever a scope is requested. For example, hardware keys using PIV device attestation provide the same X.509 certificate for each request. This design introduces two risks: (1) without state, an attacker can replay cached scope proofs; and (2) fixed values can be reused across contexts, enabling misuse outside the intended signing event.
To support SPs, DiVerify provides two mechanisms:
  • Zk-based SP: The provider is wrapped in a process that produces a zero-knowledge proof attesting that the value corresponds to a valid device fingerprint. This proof can be verified by both the DiVerify daemon and the verifier, and incorporates a nonce synchronized with other scope providers.
  • nonce-hash SP: In cases where authentication suffices and only event binding is required, the provider output is hashed together with a nonce. The provider returns a tuple (noncenonce, measurementmeasurement, sigsig(hashhash(noncenoncemeasurementmeasurement)))
Opaque Providers (OP) allow nonce injection to prevent replay attacks, but lack a reliable way to verify their identity, making them susceptible to spoofing if used in isolation. For example, this is the case of TPM ME and device fingerprinting writ large.
In order to support these, we introduce a broker/wrapper process that can be authenticated to provide opaque provider values. This is trivially done by wrapping existing scope provider software inside of a trusted execution environment.
Untrusted Providers (UT) support neither nonce injection nor provider verification, offering no protection against replay or impersonation attacks and representing the weakest security posture. We combine the strategies for OPs and SPs to support UTs.
For all the above, we introduce function prototypes to support the implementation of new scope providers. In general, for all providers other than VP, we utilize two strategies. The first is wrapped providers (i.e., wrapping the provider process within a trusted process). The second is TEE-based providers (i.e., transferring the provider process to a TEE).
6. Policy Semantics and Lifecycle
As discussed in section 5.2, verifiers rely on policies to decide when a signature is trustworthy. They must interpret evidence under consistent trust expectations that can evolve over time. This section describes how DiVerify encodes and securely updates those expectations.
6.1. DiVerify Policy Language
DiVerify policies allow software maintainers to state the trust conditions under which a signature should be accepted. Signing clients produce proofs that contain authenticated scopes and an integrity attestation of the signing tool. A verifier must determine whether this evidence satisfies the expectations expressed in the policy.
These expectations vary across settings. For example, a signer working on a personal computer may authenticate with an identity provider, present a hardware token, and show that signing occurred on a specific trusted device. When signing from a remote environment, the same signer may still prove identity or possession of a hardware key but may no longer be able to prove use of the expected device. DiVerify must support this practical flexibility. Verifiers should be able to express strict requirements when needed, as well as fallback trust paths when some contextual evidence is unavailable. Since DiVerify aggregates identity, contextual, and attestation evidence into a structured proof, verifiers require a policy language that can express such trust requirements in a deterministic and verifiable manner.
The DiVerify policy language is designed to meet these goals. First, it aligns with the structure of the DiVerify proof so that verifiers can evaluate evidence without ambiguity. Second, it is expressive enough to capture realistic signing workflows, including cases where different forms of authentication may be acceptable. Third, it is safe: a policy cannot be satisfied by combining claims from multiple signers or from inconsistent signing contexts.
To achieve this, a policy specifies constraints over three categories of evidence produced during proof construction: authenticated scopes, attestation properties of the signing environment, and a logical rule that describes how these claims must combine to satisfy the maintainer’s trust requirements.
Formally, a policy is a tuple:
$P=\langle Scope,Attest,Rule\rangle$
Scope describes what authentication claims may exist in the signature’s DiVerify proof. It is defined per signer and is structured as a collection of typed claims. For a signer $u$, the scope is modeled as:
$Scope(u)=\{\langle t_{1},v_{1}\rangle,\langle t_{2},v_{2}\rangle,\ldots,\langle t_{k},v_{k}\rangle\}$
where each $t_{i}$ denotes a scope type and $v_{i}$ is the value to be asserted by the corresponding Scope Provider.
Because a signer may possess multiple values for a given scope type (e.g., multiple trusted devices), scopes are modeled as a mapping from scope types to sets of acceptable values. For a signer $u$, $Scope(u)[t]=\{v_{1},\ldots,v_{n}\}$ is the set of values verifiers can trust. At verification time, Let $\Sigma=\{\sigma_{1},\ldots,\sigma_{n}\}$ be the set of different scopes contained in a DiVerify proof. A scope predicate $\textsf{hasScope}(t,x)$ evaluates to true if there exists a scope $\sigma\in\Sigma$ of type $t$ whose value equals $x$.
AttestAttest captures verifier-side trust requirements over the integrity of the signing environment and is modeled as a tuple
$Attest=\langle M,B,Q\rangle$
where $M$ is the expected environment measurement, $B$ is the binding value that cryptographically links the attested environment to the signing key and the authenticated scopes, and $Q$ specifies the verification procedure for TEE quotes. Given an attestation report $q$, the predicate $\textsf{validAttest}(q,\mathit{Attest})$ holds if the quote verifies under $Q$, the included measurement matches $M$, and the binding value contained in $q$ matches $B$ as derived from the signing key and authenticated scopes.
The $Rule$ defines the conditions under which a verifier considers the provided scopes and attestation evidence sufficient. It merges two concerns: (1) value-level checks ensuring a match with expected values and (2) logical composition specifying how multiple signer claims must relate to one another. A semantic constraint is that all predicates within a Rule refer to the same signer. That is, Rule expressions do not compose evidence across different signers; the verifier evaluates whether a subject signer consistently meets all required conditions. We express these as a Boolean expression over scope assertions:
$Rule=$ $\textsf{match}(u,t,x)\mid Rule_{1}\land Rule_{2}\mid Rule_{3}\lor Rule_{4}\mid\lnot Rule_{5}$
which evaluates to true when signer $u$ has a scope of type $t$ whose value set contains $x$. Complex Rules are formed by combining these predicates using the logical operators AND, OR, and NOT. All predicates in the Rule refer to the same signer; a Rule is satisfied only if there exists a single signer $u$ for which the entire Boolean expression evaluates to true.
6.2. Policy Metadata, Update, and Revocation
In addition to the semantic content of the policy which specifies the required identity claims, attestation properties, and Rule evaluation, the system must also represent how policies evolve over time and how verifiers determine whether the policy they are using is the most current and authoritative. To support secure policy evolution, we augment each policy with a lightweight layer of policy metadata. This metadata enables verifiers to authenticate the policy, detect replayed or stale versions, and reason about both semantic updates and revocation events.
A policy instance is therefore modeled as a pair $(Meta,P)$ where P is the semantic policy described in section 6.1 and Meta is a signed metadata record that governs its validity and has the form:
$Meta=\langle\textsf{version},\textsf{epoch},\textsf{issuedAt},\textsf{expiresAt},h_{P},sig_{pol},sig_{root}\rangle$
where $h_{P}$ is the hash of policy, binding the metadata to a specific policy. The version number is monotonically increasing integer that reflects semantic updates to the policy (as shown in Figure 3). Any changes that affect the meaning of the policy, such as updating trusted signer, updating the scope requirement, or altering the Rule structure, results in an increment to the version. Because such updates change the trust configuration, the metadata for a new semantic update must be jointly authorized by the long-term policy key and a higher-privileged root key. This dual signature ($sig_{pol},sig_{root}$) requirement ensures that no single actor can unilaterally introduce or weaken trust conditions.
Figure 3. Policy metadata evolution. Vertical changes correspond to version updates that alter policy semantics, while horizontal changes correspond to epoch updates that only refresh metadata.
Not all policy updates modifies semantics. In practice, a policy issuer may need to refresh operational properties of the metadata, such as extending its validity interval or reissuing it to prevent replay, without changing the meaning of the policy. To support these non-semantic updates, the policy metadata incorporates an epoch: another monotonically increasing integer that associates a particular semantic version. Incrementing the epoch yields a new instance of metadata that binds to the same policy $P$ and has the same versionversion (as shown in Figure 3). Because epoch increments do not modify the policy’s trust semantics, they require only the signature of the current key.
The policy key defines the authority permitted to issue metadata for a given semantic version. For this reason, rotating the policy key itself is a semantic change as it changes that authority. DiVerify therefore treats policy key rotation as a semantic update that must be co-authorised by the root key.
The metadata timestamps $issuedAt$ and $expiresAt$ provide verifiers with freshness guarantees. A verifier accepts a policy if its metadata only if the current time lies within its validity intervals. These timestamps prevent an adversary from indefinitely replaying an older signed metadata record. Verifiers maintain, for each policy identifier, the highest (version, epoch) pair they have previously accepted. Upon receiving new metadata, the verifier checks its signature and validates the hash binding to $P$. It then compares the incoming metadata against its stored maxima. Any metadata record with a lower version or with a lower epoch for the same version is rejected as a rollback attempt. This small amount of persistent state suffices to ensure strong rollback resistance in offline and adversarial settings, while imposing negligible storage overhead. DiVerify requires significantly less state than the manifest and metadata caches routinely maintained by software verification tools, such as APT (Hertzog and Mas, 2014), DNF (Red Hat, [n.d.]), and TUF (gittuf, 4 29) clients. In deployments where such systems already manage trusted metadata, DiVerify’s policy metadata can be integrated or subsumed into their existing mechanisms for distribution and caching (section 7).
To handle signer compromise, policy error, and evolving trust requirements, the metadata layer provides a systematic mechanism for revocation. When the policy issuer intends to remove a previously trusted element from the policy, he publishes a new semantic version that encodes this change. Such modification alters the trust semantics and therefore requires co-authorization by the root key. Revocation may also occur implicitly such that if a new semantic version excludes or tightens conditions that earlier versions permitted, then verifiers interpreting the latest version will reject proofs that depend on the superseded trust assumptions. Because verifiers enforce monotonicity over version numbers, any trust configuration that has been superseded cannot be reinstated by replaying older metadata.
7. DiVerify in Existing Systems
(a) Legacy-Compatible Model
(b) Trusted Authentication Model
Figure 4. Two integration models for using DiVerify with existing signing systems, in addition to the Core Model ( Figure 2). (a) Legacy-Compatible (no client attestation). (b) Trusted Authentication (with client attestation). The latter protects against compromised clients.
While DiVerify defines a core trust model for identity and client verification, its architecture supports multiple realizations with varying strength of guarantees. This flexibility enables incremental adoption: existing systems can integrate DiVerify using available capabilities and progressively enforce stronger trust requirements as they become ready. In this section, we describe two DiVerify instantiations that illustrate how the same underlying mechanisms integrate with existing signing systems operating under different operational constraints.
7.1. Legacy-Compatible
The legacy-compatible variant enables stronger security guarantees in existing signing workflows with minimal disruption, at the cost of a weaker posture against client compromise.
Current solutions typically rely on a single authentication mechanism: an OIDC identity token. In this variant as shown in Figure 4a, DiVerify operates without a secure daemon. Instead, the client is trusted to honestly retrieve and present authentication claims, and the DiVerify proof becomes a collection of such claims, embedded directly into an existing root of trust (e.g., a signing certificate).
For example, in systems like Sigstore, which issues ephemeral certificates tied to OIDC identities and signing keys, the certificate could be extended to include a DiVerify proof that contains additional scopes. Similarly, in OpenPubkey, which binds signing keys to federated identity claims, the OIDC scope becomes one component in a more expressive interpretations of signer legitimacy.
Even in this relaxed model, software maintainers can define DiVerify policies that go beyond a single identity source, and constraint for such identities to sign. However, this setting sacrifices client-side integrity guarantees. Because there is no trusted daemon or attestation, the system cannot validate client integrity. A compromised client could spoof or replay scopes. While a certificate authority may verify the authenticity of claims against the issuing scope providers, this only holds for secure and verifiable providers that are able to authenticate themselves.
7.2. Trusted Authentication
The Trusted Authentication variant (depicted in  Figure 4b) is similar to the core DiVerify model but differs in when the daemon’s integrity is verified and who performs the verification.
As in the core model, the daemon retrieves scopes from scope providers. However, instead of proceeding to signing after collecting and validating scopes, the daemon defers signing until an external verifier confirms its integrity. This additional check helps prevent a compromised daemon from producing valid signatures.
A common choice for this external verifier in existing systems is a Certificate Authority, which can embed the daemon’s attestation into a signing certificate. This approach aligns with existing infrastructures such as those described in (Newman et al., 2022), where CAs verify the identity of a signer before issuing a certificate. However, the external verifier does not need to be limited to traditional CAs. Alternatives include federated identity systems, policy engines, or custom services that can enforce integrity checks before authorizing a signing operation.
By requiring the daemon’s integrity to be confirmed prior to signing, this variant provides a proactive safeguard against signature generation by a compromised daemon. However, this setting still relies on a trusted CA, which introduces a single point of failure.
8. Security Analysis
[…]
Compromise Without With DiVerify
Type Diverify
Legacy Trusted Core
Compat. Auth.
[…]
Client $Circle \CIRCLE \CIRCLE$
[…]
Certificate Authority $Circle \CIRCLE$
We consider three attack scenarios (Table 1) in identity-based signing, derived from our threat model (section 3.2) and show how our core, legacy-compatible (section 7.1) and trusted authentication (section 7.2) DiVerify variants protect against them.
Types I & II (Scope Provider Compromise)
If a scope provider were to be compromised, an attacker can issue fraudulent scopes to impersonate a legitimate signer.
Under this assumption, even if an attacker fraudulently obtains one valid identity token, they still lack the the ability to provide further scopes (e.g., they cannot fake the TEE-based attestation from the signing client). Because at least one of the multiple checks will fail, the attacker cannot produce a complete DiVerify proof to sign malicious code. Hence, a single compromised provider is not enough to break the chain of trust, thus providing security goal S1 and S2 guarantees.
Type III (Client Compromise)
In this scenario, we identify the following cases:
  1. (1)
    Bypass authentication: A compromised client may skip signer authentication and spoof scopes, producing a DiVerify proof that enables unauthorized signing.
  2. (2)
    Substitution signing: Since the client controls the ephemeral signing key, it can misuse it to sign other software, yielding valid but malicious signatures.
Naively trusting such signatures allows attackers to distribute compromised software.
DiVerify’s resilience depends on the operating mode. In Legacy-Compatible, the client is trusted to authenticate signers, but without a TEE its behavior cannot be checked. A compromised client can spoof or omit verification steps, leaving this mode vulnerable.
In contrast, Core mode and Trusted Authentication mode delegate verification to a daemon running inside a TEE, while the untrusted client only forwards scope requests. Any modification of the daemon or its execution environment is detectable through attestation. Core mode validates the signing context at verification time, while Trusted Authentication mode validates it at certificate issuance. In both cases, signatures produced by a compromised client are detectable and rejected, thereby satisfying Security Goal S3.
A remaining risk is payload substitution: e.g., Alice attempts to sign package foo, but the client forwards $\textit{foo}^{\prime}$ to the daemon. Because authentication is decoupled from content, Alice has no visibility into what is actually signed. This is mitigated if (i) the daemon displays the artifact information such as the file name and its cryptographic hash directly to Alice through a trusted channel (e.g., using GPU passthrough), without relying on the compromised client or (ii) a scope provider attests to the client’s integrity using TPM ME (Intel Corporation, 2025) or IMA (Linux Integrity Subsystem, 4 21).
Certificate Authority Compromise
When a CA is used (as in Legacy-Compatible and Trusted Authentication modes), and is compromised, they can issue signing certificates without validating the client’s attestation, or may deliberately fail to flag invalid attestation. This allows a rogue client signing to appear legitimate to downstream verifiers. As a result, malicious software may be signed and distributed with a valid certificate. DiVerify mitigates this risk to varying degrees depending on the operating mode.
Trusted Authentication mode assumes trusted CA but includes the client’s attestation within the certificate. If a verifier re-validates the embedded attestation independently rather than trusting the certificate blindly, it can detect that the CA failed to perform required checks, holding CAs accountable. This attack is not applicable in Core mode, as the CA is removed from the trust chain entirely. Instead, the verifier directly validates the client’s attestation before accepting a signature, eliminating the risk of CA compromise. Legacy-Compatible mode offers no protection against this attack. Since attestations are not part of the signing process, there is no mechanism for verifying whether the CA observed valid scopes or verified the client’s integrity before issuing the certificate.
8.1. Mitigating Past Attacks with DiVerify
We illustrate how DiVerify mitigates the attacks described in section 3.2.
  1. (1)
    ESLint (Type I): The attackers used stolen maintainer credentials to publish a malicious package. With DiVerify, identity verification would have required additional contextual scopes, not just possession of the maintainer’s token. Since the attacker could not produce these scopes, their attempt to propagate the modified package via npm would have failed.
  2. (2)
    Midnight Blizzard (Type II): The attackers leveraged OAuth tokens obtained through a compromised identity provider. If the relying services had required DiVerify’s contextual scopes beyond simply accepting valid OAuth tokens, the attacker would have been unable to authenticate through the other scope providers, and the attack would have been prevented.
  3. (3)
    Codecov (Type III): The tool was silently modified, compromising CI environments of users who fetched it. With DiVerify, the package maintainers would have published a policy defining the tool’s known trusted state. At runtime, the tool would generate an attestation, enabling verification against this trusted state. Since the modified tool would have deviated, users could have detected and flagged the compromise.
Existing systems base trust solely on possession of valid credentials or tokens, leaving verifiers blind to compromised identities, providers, or tools. DiVerify prevents these attacks by exposing diverse, verifiable security context to verifiers and enforcing policies over that context, enabling misuse to be detected and rejected at verification time rather than after an attack.
9. Implementation and Evaluation
We now assess DiVerify’s practicality by measuring signing/verification times and storage overhead. Given that our implementation is built on existing signing tooling, the only costs for adoption are those of storing and executing new processes and tools within an ecosystem.
9.1. Implementation
Table 2. Implemented Scope Providers (see section 5.3 for type). Location indicates where provider runs. Size is lines of code via sloc with default parameters.
Provider Type Location Size
OIDC (Conformance, 0 12) VP Remote 46
Device Fingerprint SP Local 22
Security Key (Yubico, [n.d.]) SP Local 28
Table 3. Average Signing & Verification Overhead Introduced by DiVerify (in milliseconds). L1–L3 indicate trust levels determining the signer scopes required during signing.
Legacy Compat. (ms) Trusted Auth. (ms) Core (ms)
L1 L2 L3 L1 L2 L3 L1 L2 L3
Signing Signing Time 50.93 81.53 76.59 329.66 328.00 323.79 73.73 68.59 86.98
Quote Gen. 22.86 26.00 25.60 25.27 24.40 25.60
Fulcio 38.40 39.15 32.99 253.06 254.48 249.06
Verification Verify Time 1483.53 1357.36 1403.96 1336.44 1397.01 1381.50 1668.88 1604.10 1659.33
Quote Verif. 216.01 187.04 211.34
We build upon the widely utilized signing provide library Securesystemslib (Lab, [n.d.]) used by tools like TUF(gittuf, 4 29), in-toto (Torres-Arias et al., 2019), and Sigstore (Newman et al., 2022). Securesystemslib provides core identity-based signing and verification abstractions, which allowed us to prototype DiVerify by augmenting existing signing and verification interfaces. We implement DiVerify by making four major changes:
First, we implemented a DiVerify-aware signing to support both signing client and verifier. The client is responsible for retrieving signer scopes, signing and generating remote attestation (section 5.1); on the verifier side, it enforces scope checks against embedded proofs and package policy (section 5.2). We also implemented separate daemon logic to support operational modes where scope retrieval and signing are delegated to a dedicated background service. Second, we adapted the client to support the models described in section 7. We adapted this client along other identity-based software signing tools such as sigstore-python (Sigstore Project, [n.d.]) with similar arguments and parameters to ensure user familiarity.
Third, we utilized existing authentication protocols to provide user scopes. We use the token beacon (Conformance, 0 12) to request identity tokens from a GitHub OIDC provider. We implemented two static scope providers: a custom one for measuring device fingerprint, and another for attesting(Yubico, [n.d.]) the signer’s security key. See Table 2. We advertise supported scopes using configurable trust levels (shown in Listing LABEL:lst:diverify-config-structure) that shows progressively stronger authentication and context signals. These levels range from OIDC-based identity assertions (Level 1), to the inclusion of a hardware-backed security key (Level 2), and to device-bound identity assertions (Level 3).
Finally, we customized Sigstore’s Fulcio certificate authority (Sigstore, [n.d.]) to incorporate DiVerify proof verification prior to certificate issuance. This required modifying 201 lines of code. Verified proofs are embedded in the issued certificate.
To support remote attestation, we deployed a local Provisioning Certificate Caching Service (Corporation, [n.d.]a), which supplies quote generation and verification services with Intel SGX collateral: Provisioning Certification Key certificates, Certificate Revocation List, Trusted Computing Base info, and Quoting Enclave identity data. For quote verification, we use Intel’s SGX DCAP Quote Verification tool (Corporation, [n.d.]b).
Experiment Setup
We ran our experiments using Docker containers on a 4.0 GHz Intel Xeon Platinum 8580 machine with 499GB of RAM and SGX support, reflecting a realistic industry-grade setup for scalable signing tasks (Kalu et al., 2025b). TEE components were built using Gramine (Contributors, [n.d.]), which automatically handles the loading and execution of apps within the enclave.
9.2. Performance Evaluation
We evaluated DiVerify’s performance across the three modes, legacy-compatible, Trusted Authentication, and Core.
We evaluated different policies to mode mapping with constraints to mitigate the attack vectors in our threat model. To evaluate DiVerify’s performance overhead, we averaged over 10 iterations of artifact signing and verification for each combination of the implemented modes and trust levels.
9.2.1. Performance Overhead
Table 3 presents breakdown of time spent across the signing and verification phases. The primary contributor to signing latency is the Fulcio CA. This is evident in Trusted Authentication Mode, where the CA verifies client integrity attestation, leading to longer signing times compared to the Legacy, which skips attestation.
Signing using Core mode offers stronger security guarantees with minimal overhead. It provides less overhead (13ms) compared to the legacy-compatible variant because all operations are handled within the daemon, avoiding round trips to an external CA. Still, both signing modes provide acceptable overhead since they complete in under 100ms. However, verification in Core is more expensive due to additional attestation verification performed by the verifier, resulting to about 11% average latency increase compared to Legacy-Compatible.
Despite the added security guarantees, DiVerify’s performance remains comparable to state-of-the-art benchmarks (Newman et al., 2022). Core mode incurs slightly higher overhead than the other modes due to the additional verifier-side effort in validating the client’s integrity attestation. Nevertheless, the overhead remains within acceptable bounds and on par with current signing and verification tools.
9.2.2. Storage Overhead
We measured the size of DiVerify proofs across different operational modes, ranging from 0.4 KB (minimal identity attributes) to 6.7 KB (with full scope and attestation). This size remains fixed for deployments with consistent scope requirements. To understand the practical impact, we measured the size of the top 20 most downloaded PyPI packages in the past month (PyPI Stats, [n.d.]) and identified an average package size of 1.68 MB. DiVerify adds at most 0.39% overhead to this average size — a minimal cost relative to the package size. While the relative impact decreases for larger packages, the added size remains modest even for the smaller ones.
10. Discussion
In the interest of practicality, there are further considerations and extensions that lie outside of our system as presented thus far.
10.1. DiVerify & Signer Privacy
Our work thus far has focused on security and deployment objectives, not on anonymity or privacy for the signer. DiVerify, as described, ties a lot of identifiable information into the signature bundle, which might concern privacy-conscious signers or organizations. For instance, the certificate will reveal exactly which identity, and scope providers were used. To address this privacy concern, DiVerify can integrate Speranza (Merrill et al., 2023). Speranza relies on a chain-of-custody proof (i.e., identity co-commitments) from a repository and an IdP. Integrating DiVerify with Speranza would require to adapt scope providers to provide a portion of an identity CoCommitment. Consequently, while the set of providers remains visible, it becomes impossible to decipher the user who authenticated with them. This effectively provides privacy for signers, ensuring their anonymity when signing.
10.2. Applicability/Deployment Considerations
DiVerify’s design emphasizes backward compatibility. It builds on standard authentication flows and adds contextual identity “scopes” to signatures, allowing it to plug into existing code signing systems without replacing any infrastructure. Because of this, organizations can introduce DiVerify alongside current workflows; existing verification processes still work (even if they ignore the extra scope data). This means teams can roll out DiVerify incrementally – for example, starting with basic identity checks and gradually adding more stringent scope requirements as they become ready. In other words, current flows are effectively a rather constrained instantiation of DiVerify, as far as existing clients are aware.
Integrating DiVerify requires little changes to existing tooling. Our implementation illustrates this. The integration only required 2007 lines of code for the client side signing and verification tooling, and 201 lines of code increase for Fulcio. These changes imply maintainers can add DiVerify with very little development effort.
Because DiVerify leverages existing software signing architecture, maintainers need not change their environments, but rather augment them. Another consideration is the choice of scopes for a given context. Our approach is flexible across the claim types identified in section 2.2. In automated build systems, DiVerify could be used to ensure that build and signing servers themselves authenticate via scopes (like machine certificates or runner attestations), and that triggers for signing (like a GitHub Actions workflow run) carry a scope proving an authorized user initiated them.
As with multi-factor authentication systems (De Cristofaro et al., 2014), DiVerify does not assume a single notion of authentication effort. Strong security improvements generally require additional, non-substituable evidence, and therefore introduce trade-offs that vary across deployments. DiVerify is designed to make these trade-offs manageable by allowing maintainers to choose which combination of identity scopes are required for acceptance.
11. Related Work
This work relies on existing works from access control mechanisms, as well as distributed identity mechanisms. We contextualize DiVerify in these techniques, focusing on their applications to software supply chains.
Access Control Mechanisms
Role-Based Access Control (RBAC) has long been foundational in managing permissions (Sandhu et al., 1996), with widespread adoption in general-purpose systems (Microsoft Azure, 4 21; Microsoft, 4 29), cloud platforms (kubernetes, 4 29; Ansible Project, 2024), and social coding environments (GitHub, 2024; GitLab, 2024). In supply chain contexts, Kuppusami et al. (Kuppusamy et al., 2016; gittuf, 4 29) propose repository-specific models for permission management, leveraging abstractions like branches and commits. DiVerify differs by mapping permissions across heterogeneous domains without sacrificing security guarantees, particularly in the identity-based software signing context.
Several studies have also identified limitations in RBAC implementations for supply chain security (Moore et al., 2023; Koishybayev et al., 2022), focusing on repository metadata and CI platforms. These insights inform DiVerify’s design, especially in modeling semantics across diverse systems.
Verifiable Identities and Credentials
Similarly, the application of identity for different contexts has surfaced in general identity work. Perhaps one of the richest lines of work in this regard is that of Verifiable Credentials. Prior works show how credential holders can selectively and verifiably disclose properties they possess (Johnson et al., 2023; Lux et al., 2020; Mukta et al., 2020; Lim et al., 2021; OpenID Foundation, nd) DiVerify provides a similar goal by building upon existing and widely deployed technologies (such as OpenID Connect). Properties such as the ability to commit to a repository or holding ownership of a project can be achieved by off-the-shelf algorithms and tools.
Similarly, applications of distributed identity frameworks (i.e., DiD (World Wide Web Consortium, 2022; bcGove, 2023; walt.id, 2023)) can be used in lieu of identity providers. This has been considered in standards such as SCITT as a possible building block for a global software supply chain ledger (IETF, 2022). Were these standards to become widespread, DiVerify can leverage them. However, it is necessary to apply the policy logic to merge various providers described in this paper (section 5.3) to ensure proper capabilities are applied.
Threshold-based Authentication
Threshold-based authentication schemes, which distribute trust across multiple identity providers (Ito et al., 2017) using secret sharing (Shamir, 1979), have been proposed for enhancing resilience and privacy in federated identity systems. These approaches align with DiVerify’s goal of federated identity verification but typically rely on a trusted dealer. DiVerify avoids this by ensuring no single party can subvert trust, maintaining a decentralized model
12. Conclusion
We presented DiVerify, a novel approach to strengthening software signing systems against threats from both identity provider and client compromise, two key problems in current signing architectures. DiVerify addresses these by collecting diverse, verifiable trust signals for both signers and clients, enabling quick detection of malicious signing and improving the resilience of software signing as a security measure.
Our prototype implementation demonstrates that DiVerify is both practical and effective, enabling proactive threat mitigation with minimal performance overhead. DiVerify establishes a security baseline that signing tools can adopt to provide robust code-signing guarantees. We leave integration with build systems and measuring usability as future work.
Appendix
The appendices contain the following material:
  • Appendix A: An extended version of the assumptions we made.
  • Appendix B: Sample DiVerify Artifact Structure
  • Appendix C: Sample policy used in our analysis, with signer scope mapped to the expected claims, and the constraint the signature must meet to be trusted.
Appendix A Extended Assumptions
Beyond the attacker models described in section 3.2, we make standard assumptions from previous literature (Cappos et al., 2008; Newman et al., 2022). First, attackers are able to interpose communications, but standard mechanisms to prevent person-in-the-middle attacks (e.g., TLS) are in place and all parties utilize these. Similarly, attackers are able to replay communication messages to attempt to impersonate roles in the system.
However, an attacker is not able to:
  1. (1)
    Break standard cryptographic algorithms (e.g., sha256, ed25519). An adversary capable of breaking standard cryptographic algorithms is out of scope.
  2. (2)
    Overcome kernel-level or TEE-based system measurement and isolation techniques. While attacking TEEs is a thriving area of offensive security research, we assume that these systems are functioning correctly.
  3. (3)
    Compromise a collection of IdPs. While a sophisticated attacker could compromise large swaths of identity infrastructure, this is rarely the case in practice due to the practical difficulty of breaching multiple independent scope providers at once, especially when operated by different organizations with diverse infrastructures and security postures.
Lastly, we assume that the Signer is legitimate (i.e., behaves honestly and does not intend to signing of untrusted code). We also assume that the verifier behaves honestly and correctly verifies signatures and associated identity information according to the expected protocol.
Appendix B Other DiVerify Artifact Samples
Listing LABEL:lst:diverify-config-structure defines a sample trust level configuration used in DiVerify, specifying the scopes required for a signer at each level. Level 1 provides identity claims, level 2 adds possession claims, and level 3 includes contextual claims in addition to level 2.
Listing LABEL:lst:proof-structure shows a sample structure of the DiVerify proof, showing the sample collected scopes that represent claims about both the signer and the client.
1{
2 "Level": 1,
3 "identity": {
4 "oidc": true
5 },
6 "Level": 2,
7 "identity": {
8 "oidc": true,
9 "security_key": true
10 },
11 "Level": 3,
12 "identity": {
13 "oidc": true,
14 "security_key": true,
15 "device_fingerprint": true
16 }
17}
Listing 1: Sample Trust Level
1{
2"config_version": "1.2",
3"trust_level": 3,
4"identity": {
5"oidc": {
6 "identity": "...",
7 "provider": "...",
8 "token_hash": "...",
9 },
10"device_fingerprint": "...",
11"security_key": {
12 "slot9a_public_key": "...",
13 "slotf9_attestation_cert": "..."
14 },
15 },
16"signing_key": "..."
17"remote_attestation": {
18 "enclave_quote": "..."
19 }
20}
Listing 2: Sample DiVerify Proof
Appendix C Test DiVerify Policies
Figure 5 presents a sample policy used to evaluate whether to trust a package signature.
Figure 5. An example DiVerify policy specifying expected claims. The final field is a Boolean rule constraint under which a package signature is to be trusted. This policy is defined by the software producer and enforced by the verifier, ensuring that only signatures that meet this constraint are accepted, which strengthens the overall trust in the signing process.